### 17 Mar 2026 v1

# Nearly-polynomial inverse theorem for the $U^{d}$ norm in degree $d+1$

Abstract
We prove a nearly polynomial inverse theorem for the Gowers $U^{d}$ norm, over finite fields of non-small characteristic, for polynomials of degree $d+1$. The case of degree $d$ was very recently settled by Milićević and Rand¯elović with a fully polynomial bound. We moreover provide a nearly polynomial inverse theorem for homogeneous polynomials of any degree smaller than $2d$.
1 Introduction
Let $\mathbb{F}=\mathbb{F}_{p}$ be any prime finite field. The discrete derivative of a function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ in direction $v\in\mathbb{F}^{n}$ at the point $x\in\mathbb{F}^{n}$ is $\Delta_{v}f(x)=f(x+v)-f(x)$. More generally, the (discrete) derivative of order $d$, in directions $\mathbf{v}=(v^{(1)},\ldots,v^{(d)})\in(\mathbb{F}^{n})^{d}$ at the point $x\in\mathbb{F}^{n}$, is
$\Delta^{d}_{\mathbf{v}}f(x):=\Delta_{v^{(d)}}\cdots\Delta_{v^{(1)}}f(x).$
A function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ is a polynomial of degree below $d$, denoted $f\in\operatorname{Poly}_{<d}$, if and only if $\Delta^{d}f=0$ (that is, $\Delta^{d}_{\mathbf{v}}f(x)=0$ for every $(x,\mathbf{v})\in(\mathbb{F}^{n})^{d+1}$).
An important (qualitative) result about polynomials over $\mathbb{F}$, of characteristic $\operatorname{ch}(\mathbb{F})>d$, is that the derivative characterization is robust: $f$ is correlated with $\operatorname{Poly}_{<d}$ if and only if $\Delta^{d}f$ is biased. This lies at the core of the theory of Gowers norms [Gow01, Tao12, HHL19], and is usually formulated using exponential sums: [^1] Which measure the distance of a distribution to uniformity; see e.g. Claim 33 in [BV10]. for any function $f\colon\mathbb{F}^{m}\to\mathbb{F}$, one denotes
$\operatorname{bias}(f)=|\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{m}}e(f(x))|\quad\text{and}\quad\operatorname{cor}(f,P)=\operatorname{bias}(f-P)$
where $e(y):=e^{2\pi iy/p}\colon\mathbb{F}\to\mathbb{C}$ and $\operatorname*{\mathbb{E}}_{x\in A}:=\frac{1}{|A|}\sum_{x\in A}$. We note that $\operatorname{bias}(f)\in[0,1]$, [^2] For example, if $\deg(f)<d$ then $\operatorname{bias}(\Delta^{d}f)=1$, while if $f$ is uniformly random then $\operatorname{bias}(\Delta^{d}f)=o(1)$. and that, explicitly, $\operatorname{bias}(\Delta^{d}f):=\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n},\mathbf{v}\in(\mathbb{F}^{n})^{d}}e(\Delta^{d}_{\mathbf{v}}f(x))$. The forward direction of this characterization—correlation implies bias—is not hard to prove, even quantitatively (e.g., using the Cauchy-Schwarz-Gowers inequality):
$\operatorname{bias}(\Delta^{d}f)\geq\operatorname{cor}_{<d}(f)^{2^{d}}$ (1)
where $\operatorname{cor}_{<d}(f):=\max_{\deg(P)<d}\operatorname{cor}(f,P)$.
The inverse direction is much harder. For $d=2$, the question falls in the domain of Fourier analysis, as $\operatorname{bias}(\Delta^{2}f)=\sum_{\zeta}|\hat{f}(\zeta)|^{4}$, and Parseval’s identity implies $\operatorname{cor}_{<2}(f)\geq\operatorname{bias}(\Delta^{2}f)^{1/2}$. However, standard Fourier analysis is insufficient for general $d$, and the inverse direction is only known to hold with an extremely weak bound [GM20]. The inverse direction is often stated in terms of the Gowers (uniformity) norms of a complex-valued function. For $g\colon\mathbb{F}^{n}\to\mathbb{C}$, the $U^{d}$-norm is $\lVert{g}\rVert_{U^{d}}=(\operatorname*{\mathbb{E}}_{(x,\mathbf{v})\in(\mathbb{F}^{n})^{d+1}}{\Delta^{*}}_{v^{(d)}}\cdots{\Delta^{*}}_{v^{(1)}}g(x))^{1/2^{d}}$, with $\Delta^{*}_{v}g(x)=g(x+v)\overline{g(x)}$, and the weak Gowers norm is $\lVert{g}\rVert_{u^{d}}=\max_{\deg(P)<d}\big|\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n}}g(x)e(-P(x))\big|$.
Theorem 1.1 (Gowers inverse over finite fields [GT09, BTZ10, TZ10, TZ12]).
For any $g\colon\mathbb{F}^{n}\to\mathbb{C}$ with $\|g\|_{\infty}\leq 1$,
$\lVert{g}\rVert_{u^{d}}\geq\Phi(\|g\|_{U^{d}})$
for some function $\Phi=\Phi_{d,\mathbb{F}}>0$.
In the representative case where $g=e(f)$ with $f\colon\mathbb{F}^{n}\to\mathbb{F}$ (e.g., Lemma 1.2 in [TZ10]), the Gowers norm is simply $\lVert{g}\rVert_{U^{d}}=\operatorname{bias}(\Delta^{d}f)^{1/2^{d}}$, so Theorem 1.1 is a qualitative inverse to (1):
$\exists\Phi>0\colon\operatorname{cor}_{<d}(f)\geq\Phi(\operatorname{bias}(\Delta^{d}f)).$ (2)
We note that if the field has small characteristic $\operatorname{ch}(\mathbb{F})<d-1$, the maximum in Theorem 1.1, over $P$ with $\deg(P)<d$, must also range over so-called non-classical polynomials of degree below $d$ ([BS+22, GT09, LMS11]), which are functions $\mathbb{F}^{n}\to\mathbb{R}/\mathbb{Z}$ that involve division by powers of $\operatorname{ch}(\mathbb{F})$. In this paper we restrict ourselves to the high characteristic case (and so, to classical polynomials).
The Polynomial Gowers Inverse conjecture posits that, in fact, the relationship between bias and correlation is about as tight as it could be; namely, that $\Phi(x)$ in Theorem 1.1 can be as large as a power of $x$ for any fixed $d$: $\Phi(x)\geq x^{O_{d}(1)}$ (this matches the easy direction $\Phi(x)\leq x^{2^{d}}$ in (1) up to the constant). It is perhaps the central conjecture of higher-order Fourier analysis, over finite fields and over the integers (e.g. [LSS24]), with applications in countless fields (see, e.g., [HHL19, Tao12, TZ10] and the references within).
While a recent breakthrough in [GG+24, GG+25] was achieved in the case $d=3$, the Polynomial Gowers Inverse conjecture for any $d>3$ remains wide open.
1.1 Gowers inverse in low degree
[…]
In this paper we obtain an almost-polynomial bound for $\operatorname{GI}_{d}(d+1)$, the next open case. With a slight abuse of notation, we write $x^{\tilde{O}_{d}(1)}$ for $\exp(-O_{d}(\log(1/x)^{1+o(1)}))$. (This is even stronger than a quasi-polynomial bound, which has a constant $O_{d}(1)$ at the top instead of 1+o(1)1+o(1).) [^4] More precisely, we get $\log(1/x)\cdot\log(\log(1/x)+1)$ in the exponent. (In fact, we even get a bound that improves as $|\mathbb{F}|$ increases, of the rough shape $x^{O_{d}(\log_{|\mathbb{F}|}(\log_{|\mathbb{F}|}(1/x)))}$, which for $|\mathbb{F}|\gg\log(1/x)$ is truly polynomial $x^{O_{d}(1)}$; see [MZ22].)
Theorem 1.
Let $d\geq 1$. For every $f\colon\mathbb{F}^{n}\to\mathbb{F}$ with $\deg(f)\leq d+1$, where $d+1<\operatorname{ch}(\mathbb{F})$,
$\operatorname{cor}_{<d}(f)\geq\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}.$
Put differently, $\lVert{g}\rVert_{u^{d}}\geq\|g\|_{U^{d}}^{\tilde{O}_{d}(1)}$ for any polynomial phase $g\colon\mathbb{F}^{n}\to\mathbb{C}$ of degree at most $d+1$ (that is, $g=e(f)$ with $\deg(f)\leq d+1$).
The main difficulty in the proof of Theorem 1 stems from the non-homogeneity of $f$. Standard proofs of the degree-$d$ case proceed by finding a decomposition of $f\in\operatorname{Poly}_{d}$ into lower-degree polynomials of length bounded in terms of $\operatorname{bias}(\Delta^{d}f)$, which is then used to obtain a significant correlation with a lower-degree polynomial. However, when $\deg(f)=d+1$, one obtains in this way only a decomposition of the top-degree part of $f$, and indeed, it is entirely possible for the degree-$d$ part $f_{d}$ of $f$ to not have any short decompositions despite having large $\operatorname{bias}(\Delta^{d}f)$ (e.g., $f=(x_{1}+1)h$ for a random $h$, see Remark 6.2). Furthermore, even if one were to successfully obtain a short decomposition of $f$ in terms of polynomials of degrees below $\deg(f)$, it is not clear how to use this to obtain a correlation with a polynomial of degree as small as $d-1=\deg(f)-2$; especially given that, unlike in the degree-$d$ case, one cannot guarantee a short decomposition with homogeneous factors. [^5] Concretely, the degree-$1$ factors in a decomposition of $f$ might not have a common zero (as they need not be homogeneous; e.g., $x_{1}$ and $x_{1}+1$), and so there might not even be a large subspace, or affine subspace, where the degree-$1$ factors can be discarded (which would have meant that all factors have degree at most $\deg(f)-2$).
In the homogeneous setting, we can do more. Namely, we obtain almost-polynomial bounds for homogeneous polynomials of degree up to $2d$.
Theorem 2.
Let $d\geq 1$. For homogeneous $f\colon\mathbb{F}^{n}\to\mathbb{F}$ with $\deg(f)<2d$, where $\deg(f)<\operatorname{ch}(\mathbb{F})$,
$\operatorname{cor}_{<d}(f)\geq\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}$
The proof of Theorem 2 relies on a new, general correlation lemma. Unlike standard correlation lemmas in the literature (e.g. [Jan19], proof of Theorem 1.6, or [GT09, CM23] and many more) which require a full decomposition of the given polynomial, ours requires a more relaxed condition, closely related to membership in an appropriate ideal.
Furthermore, as another application of our new correlation lemma, we show in Section 3.1 how it can be used to provide an alternative proof of a polynomial bound for $\operatorname{GI}_{d}(d)$ from a weak version of the partition-versus-analytic rank dichotomy [Mil25].
2 Preliminaries and proof outline
2.1 Preliminaries
Henceforth, $\mathbb{F}$ is any finite field. We write $\operatorname{\mathfrak{bias}}$ (in Fraktur letters) for the bias as a complex number: for any function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ over a finite field $\mathbb{F}$, denote
$\operatorname{\mathfrak{bias}}(f)=\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n}}\chi(f(x))\in\mathbb{C}\quad\text{ and }\quad\operatorname{bias}(f)=|\operatorname{\mathfrak{bias}}(f)|\in[0,1]$
where $\chi$ is any nontrivial additive character of $\mathbb{F}$ that we henceforth fix (for example, for $\mathbb{F}=\mathbb{F}_{p}$ prime, we can take $\chi(y)=e^{2\pi iy/p}$). Note $\operatorname{cor}(f,P)=\operatorname{bias}(f-P)\in[0,1]$. We use bold letters, such as $\mathbf{A}=(A_{1},\ldots,A_{m})$, to emphasize that the object in question is a vector/tuple. We denote the indicator function for a set $S$ by $\mathbf{1}_{S}(x)=\begin{cases}1&x\in S\\ 0&x\notin S\end{cases}$.
A form is a homogeneous polynomial. We denote by $\operatorname{Form}_{k}(\mathbb{F})$ the set of forms of degree $k$, over the field $\mathbb{F}$, in any (finite) number of variables, or simply $\operatorname{Form}_{k}$ when $\mathbb{F}$ is immaterial. Similarly, we denote by $\operatorname{Poly}_{k}(\mathbb{F})$ the set of polynomials of degree at most $k$ over $\mathbb{F}$ is any number of variables. Note that any $f\in\operatorname{Poly}_{k}$ is the sum of forms, namely, the homogeneous components of $f$ of each degree: $f=\sum_{i=0}^{k}f_{i}$ with $f_{i}\in\operatorname{Form}_{i}$. We also denote $f_{<d}:=\sum_{i=0}^{d-1}f_{i}$.
The rank of a form (also called strength or Schmidt rank), denoted by $\operatorname{rk}(f)$, is the minimal $r$ for which $f$ is the sum of reducible forms:
$f=\sum_{i=1}^{r}g_{i}h_{i}$
for some forms $g_{i}$ and $h_{i}$ of degree below $d$. A form is multilinear if it is linear in each of its variable vectors; more concretely, a multilinear form of degree $d$, or a $d$-linear form for short, is of the form $T=T(x^{(1)},\ldots,x^{(d)}):(\mathbb{F}^{n})^{d}\to\mathbb{F}$ and is linear in each of $x^{(1)},\ldots,x^{(d)}$. Equivalently, any monomial in the support of $T$ has one variable from each variable vector. The partition rank [Nas20] of a $d$-linear form $T$, denoted by $\operatorname{PR}(T)$, is the minimal $r$ for which $T$ is the sum of reducible $d$-linear forms:
$T=\sum_{i=1}^{r}R_{i}Q_{i}$
for some $R_{i}=R_{i}(x_{I_{i}})$ and $Q_{i}=Q_{i}(x_{I_{i}^{c}})$ with $\emptyset\subsetneq I_{i}\subsetneq[d]$, where for $x\in(\mathbb{F}^{n})^{d}$ we write $x_{I}$ for the restriction of $x$ to the coordinates corresponding to $I\subseteq[d]$. A related notion to the partition rank is that of the analytic rank [GW11] over a finite field $\mathbb{F}$, denoted by $\operatorname{AR}(T)$:
$\operatorname{AR}(T):=-\log_{|\mathbb{F}|}(\operatorname{bias}(T)).$
We note that for a multilinear $T$, we have $\operatorname{bias}(T)=\mathbb{P}(T=0)-\mathbb{P}(T=c)$ for any nonzero $c\in\mathbb{F}$.
The polarization of a form $g\in\operatorname{Form}_{k}$, which we denote by $\overline{g}$, is the unique (degree-$k$) multilinear form satisfying $\overline{g}(x,\ldots,x)=k!g(x)$. (Note that if $g\colon\mathbb{F}^{n}\to\mathbb{F}$ then $\overline{g}\colon(\mathbb{F}^{n})^{k}\to\mathbb{F}$.) Equivalently, $\overline{g}$ can be obtained as $\overline{g}=\Delta^{k}g$, the discrete derivative of order $k$; that is, $\overline{g}(\mathbf{x})=\Delta^{k}_{\mathbf{x}}g$ where $\mathbf{x}=(x^{(1)},\ldots,x^{(k)})$. (See Fact 5.7 below.) Note that $\Delta^{k}_{\mathbf{x}}g$ depends only on $\mathbf{x}$, being independent of the point at which we take the discrete derivative (as $k=\deg(g)$). For example, for $g(x)=x_{1}x_{2}$, we have $\overline{g}(x,y)=x_{1}y_{2}+y_{1}x_{2}$, and for $g(x)=x_{1}^{2}$ we have $\overline{g}(x,y)=2x_{1}y_{1}$.
We denote by $\partial_{c}f$ the formal derivative of a polynomial $f$ in direction $c$, and denote by $\boldsymbol{\nabla}f$ the formal gradient of $f$, so that $\partial_{c}f(x)=c\cdot\boldsymbol{\nabla}f$. Note that the formal derivative is well defined for polynomials over any field; by linearity, it suffices to define it for monomials: for $f=\prod_{i}x_{i}^{d_{i}}$, $\frac{\partial f}{\partial x_{j}}:=d_{j}x_{j}^{d_{j}-1}\prod_{i\neq j}x^{d_{i}}$, $\boldsymbol{\nabla}f:=(\frac{\partial f}{\partial x_{i}})_{i=1}^{n}$ and $\partial_{c}f=c\cdot\boldsymbol{\nabla}f=\sum_{i=1}^{n}c_{i}\frac{\partial f}{\partial x_{i}}$.
A fundamental part of our argument relies on the best bounds known for the rank of biased homogeneous polynomials (see, e.g., Corollary 8.5 in [MZ22] [^6] One can naively extend the definition of $\operatorname{rk}$ to non-homogeneous polynomials $f\in\operatorname{Poly}_{k}$ by defining $\operatorname{rk}(f):=\operatorname{rk}(f_{k})$, but this is not useful in this paper as this does not give a decomposition of $f$, only of $f_{k}$.). We write $\tilde{O}_{d}(x):=O_{d}(x^{1+o(1)})$. [^7] More precisely, we can get $\tilde{O}_{d}(x)=O_{d}(x\log(x+1))$. (In fact, we can even get $\tilde{O}_{d}(x)=O_{d}(x(\log_{|\mathbb{F}|}(x+1)+1))$, which in particular is linear for mildly large $\mathbb{F}$).
Theorem 2.1 ([MZ22]).
For every form $g\in\operatorname{Form}_{d}$, over any finite field $\mathbb{F}$ with $\operatorname{ch}(\mathbb{F})>d$,
$\operatorname{rk}(g)\leq\tilde{O}_{d}(\operatorname{AR}(g)).$
Moreover, for every $d$-linear form $T$,
$\operatorname{PR}(T)=\tilde{O}_{d}(\operatorname{AR}(T))$
2.2 Proof outline
We start by sketching the standard proof of $\operatorname{GI}_{d}(d)$. The argument is roughly as follows. For $f$ of degree $d$, the $d$-th derivative $\Delta^{d}f$ is a degree-$d$ multilinear form, which is biased by assumption. Call that bias $\delta$. Using a partition-versus-analytic rank bound, such as Theorem 2.1, one obtains a partition-rank decomposition for $\Delta^{d}f$ of length $\log(\frac{1}{\delta})^{\tilde{O}_{d}(1)}$. By plugging in the diagonal $(x,\ldots,x)$ and dividing by $d!$, we obtain a decomposition of $f_{d}$ (the degree-$d$ part of $f$), meaning $\operatorname{rk}(f_{d})\leq\log(\frac{1}{\delta})^{\tilde{O}_{d}(1)}$. Writing $f=\sum_{i}\alpha_{i}\beta_{i}$, a standard Fourier/pigeonhole-type argument gives a correlation for $f_{d}$ of at least $\delta^{\tilde{O}_{d}(1)}$ with some member of the span of the $\alpha_{i}$ and $\beta_{i}$—which is of degree at most d1d-1—and thus a correlation for $f$ itself with such a lower-degree polynomial.
The proof of Theorem 2, for homogeneous polynomials, follows this classical recipe, except it utilizes a new correlation lemma (Lemma 3.1) to obtain a correlation with a linear combination of only the lower-degree factors from each product $\alpha_{i}\beta_{i}$ in the decomposition, which is a polynomial of degree as low as $\deg(f)/2$.
The difficulty in the non-homogeneous case of Theorem 1 is twofold. First, $\Delta^{d}f$ is no longer homogeneous of degree $d$, so a short decomposition of $\Delta^{d}f$ is not immediately available from its high bias. A significant part of this paper (Sections 4 and 5) is dedicated to the task of carefully obtaining a special decomposition for a general $f$ of degree $d+1$ which—while not being necessarily a homogeneous decomposition—can be used to obtain a correlation with a polynomial of degree at most $\deg(f)-2$. We briefly sketch now how this is done.
In Section 4, we introduce a new notion of rank for polynomials called $\operatorname{rk}^{*}$. The objective here was to find a notion of rank for which the following holds: whenever $\operatorname{rk}^{*}(f)$ is small, one can find a large correlation of $f$ with a polynomial of degree at most $\deg(f)-2$ (whereas for the usual notion of rank, one can only generally get $\deg(f)-1$). It turns out that the requirement needed is quite weak: if $f=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ is a $\operatorname{rk}$-decomposition, for it to be a $\operatorname{rk}^{*}$ decomposition, we require the degree-$1$ polynomials in the decomposition to have linearly independent linear components. This requirement guarantees that the zero-locus of these polynomials is fairly large, which is essential for us to obtain significant correlation.
We now explain how the $\operatorname{rk}^{*}$-decomposition of $f$ is obtained. Write $g=f_{d+1}$ and $h=f_{d}$ for the homogeneous parts of degree $d+1$ and $d$ respectively. Proposition 5.1 provides a simple formula for $\Delta^{d}f$ for any polynomial $f$ of degree $d+1$. Proposition 6.4 shows that decomposing $g$ by the standard polarization process goes smoothly, thanks to the chain rule for bias (Lemma 6.3), which tells us that $\Delta^{d}g$} is biased. The decomposition of $h$ is a more delicate issue, since the polarization $\overline{h}$ need not be biased. However, we prove that $\overline{h}$ is biased when adding an appropriate partial derivative $\partial_{c}\overline{g}$; moreover, the sum $\overline{h}+\partial_{c}\overline{g}$ is a $d$-linear form, hence another application of Theorem 2.1 gives a decomposition to this sum. To deal with the summand $\partial_{c}\overline{g}$ we added, we use a chain rule for $\operatorname{rk}^{*}$ (Proposition 4.5).
All of the above allows us to obtain a non-homogeneous decomposition for $f$, which is nevertheless useful for us since it is a short $\operatorname{rk}^{*}$-decomposition. Using the control we have over its linear factors, Proposition 4.7 allows us then to use this special decomposition to obtain significant correlation with a polynomial of degree $\deg(f)-2\leq d-1$, as needed.
We finally remark that we cannot utilize the arguments leading to the polynomial bound for $\operatorname{GI}_{d}(d)$, in Theorem 3.4, in order to obtain a similar bound for degree $d+1$, since our argument requires more control over the factors than is afforded by that Theorem.
3 Correlation lemma, the homogeneous and the degree-$d$ cases
In this section we prove a correlation lemma which will be important in the rest of the paper. As an almost immediate corollary, we will obtain Theorem 2.
For a tuple of functions $\mathbf{A}=(A_{1},\ldots,A_{m})$, their joint zero set is $\operatorname{Z}(\mathbf{A})=\{x\,|\,\forall i\colon A_{i}(x)=0\}$. For an event $E$, we abbreviate $\operatorname{\mathfrak{bias}}(g\,|\,E)=\operatorname*{\mathbb{E}}[\chi(g)\,|\,E]$.
Lemma 3.1 (average correlation lemma).
Let $g,A_{1},\ldots,A_{m}\colon\mathbb{F}^{n}\to\mathbb{F}$ be any functions, and put $\mathbf{A}=(A_{1},\ldots,A_{m})$. Then
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\cdot\operatorname{\mathfrak{bias}}(g\,|\,\operatorname{Z}(\mathbf{A})).$
Proof.
By direct computation:
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)$ $=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\chi\Big(g(x)-\sum_{i=1}^{m}c_{i}A_{i}(x)\Big)$
$=\operatorname*{\mathbb{E}}_{x}\chi(g(x))\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\chi\Big(\sum_{i=1}^{m}c_{i}A_{i}(x)\Big)$
$=\operatorname*{\mathbb{E}}_{x}\chi(g(x))\mathbf{1}_{\operatorname{Z}(\mathbf{A})}(x)$
$=|\mathbb{F}|^{-n}\sum_{x\in\operatorname{Z}(\mathbf{A})}\chi(g(x))$
$=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\frac{1}{|\operatorname{Z}(\mathbf{A})|}\sum_{x\in\operatorname{Z}(\mathbf{A})}\chi(g(x))$
$=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\operatorname{\mathfrak{bias}}(g\,|\,\operatorname{Z}(\mathbf{A})).\qed$
We will also need the following Theorem due to Warning (for a short proof see [Asg23]).
Theorem 3.2 (Warning’s second theorem [War35]).
For all polynomials $f_{1},\ldots,f_{m}\colon\mathbb{F}^{n}\to\mathbb{F}$ over a finite field $\mathbb{F}$, if $\operatorname{Z}(f_{1},\ldots,f_{m})\neq\emptyset$ then $|\operatorname{Z}(f_{1},\ldots,f_{m})|\geq|\mathbb{F}|^{n-\sum_{i=1}^{m}\deg(f_{i})}$.
We obtain the following correlation bound, where $\operatorname{cor}_{d}(g)=\max_{\deg(P)\leq d}\operatorname{cor}(g,P)$.
Corollary 3.3.
If $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(g)$ for some $\mathbf{A}=(A_{1},\ldots,A_{m})\in\operatorname{Form}_{d}(\mathbb{F})^{m}$, then $\operatorname{cor}_{d}(g)\geq|\mathbb{F}|^{-dm}$.
Proof.
Apply Lemma 3.1 to obtain
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{bias}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)$ $\geq\Big|\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)\Big|$
$=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\cdot\operatorname{bias}(g\,|\,\operatorname{Z}(\mathbf{A}))=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\geq|\mathbb{F}|^{-dm},$
where the last inequality applies Theorem 3.2, using that $0\in\operatorname{Z}(A_{1},\ldots,A_{m})$ by homogeneity. Thus, there exists a form $P=\sum_{i=1}^{m}c_{i}A_{i}$ of degree at most $d$ such that $\operatorname{bias}(g-P)\geq|\mathbb{F}|^{-dm}$. ∎
We now prove Theorem 2.
Proof of Theorem 2.
Let $f:\mathbb{F}^{n}\to\mathbb{F}$ be a homogeneous polynomial with $k:=\deg(f)<2d$, where $\mathbb{F}=\mathbb{F}_{q}$ with $\operatorname{ch}(\mathbb{F}_{q})>k$, and put $\operatorname{bias}(\Delta^{d}f)=\delta$. First, we observe that $\operatorname{bias}(\overline{f})\geq\delta^{2^{d}}$. Indeed, using the monotonicity of Gowers norms, $\lVert{g}\rVert_{U^{k}}\geq\lVert{g}\rVert_{U^{d}}$ for $k\geq d$, and using the fact that, by definition, $\operatorname{bias}(\Delta^{t}f)=\lVert{e(f)}\rVert_{U^{t}}^{2^{t}}$ for any $t$, we have that
$\operatorname{bias}(\overline{f})=\operatorname{bias}(\Delta^{k}f)=\lVert{e(f)}\rVert^{2^{k}}_{U^{k}}\geq\lVert{e(f)}\rVert^{2^{k}}_{U^{d}}=\operatorname{bias}(\Delta^{d}f)^{2^{k-d}}\geq\delta^{2^{d}}.$
By Theorem 2.1 applied on the $k$-linear form $\overline{f}$, we obtain a partition-rank decomposition
$\overline{f}=\sum_{i=1}^{r}Q_{i}R_{i}$
with $Q_{i},R_{i}$ multilinear forms of degrees at most k1k-1, and
$r\leq\tilde{O}_{k}(\log_{p}(1/\operatorname{bias}(\overline{f})))=\tilde{O}_{d}(\log_{q}(1/\delta)).$
Assume without loss of generality that $\deg(Q_{i})\leq\deg(R_{i})$ for every $i$. Since $\deg(Q_{i})+\deg(R_{i})=k$, we have $\deg(Q_{i})\leq k/2$. Thus, $\overline{f}\in\langle Q_{1},\ldots,Q_{r}\rangle$. Plug in the diagonal: let $A_{i}(x)=Q_{i}(x,\ldots,x)$, and observe that $f(x)=\frac{1}{k!}\overline{f}(x,\ldots,x)$, using the homogeneity of $f$ and $\operatorname{ch}(\mathbb{F})>k$. It follows that $f\in\langle A_{1},\ldots,A_{r}\rangle$ with each $A_{i}$ homogeneous of degree at most $k/2<d$, by assumption. We are now done by Corollary 3.3:
$\operatorname{cor}_{<d}(f)\geq q^{-dr}\geq p^{-\tilde{O}_{d}(\log_{q}(1/\delta))}=\delta^{\tilde{O}_{d}(1)}.\qed$
3.1 A remark on the case of degree-$d$ polynomials
In [Mil25], Milićević  proved a “weak structural result for biased multilinear forms” (see Theorem 3.4 below). In this section we show that it easily implies the degree-$d$ case of the Polynomial Gowers Inverse conjecture, using our correlation lemma, Lemma 3.1. For comparison, the degree-$d$ case was proved in [MR25] using an approximation result (Theorem 2 in that paper) which roughly states that the phase of any multilinear form of large bias is close in the $L^{2}$-norm to a linear combination of phases of lower-degree forms.
We first state the aforementioned weak structural result. Note that if a multilinear form $T\colon(\mathbb{F}^{n})^{d}\to\mathbb{F}$ is of degree $k$ then $T=T(x^{(i_{1})},\ldots,x^{(i_{k})})$ depends on some $k$ of the $d$ variable vectors $x^{(1)},\ldots,x^{(d)}$ (that is, $T\colon(\mathbb{F}^{n})^{I}\to\mathbb{F}$ for some subset $I\subseteq[d]$ with $|I|=k$).
Call a subset $V\subseteq(\mathbb{F}^{n})^{d}$ a multilinear variety if $V=\operatorname{Z}(T_{1},\ldots,T_{m})$ is the zero set of a collection of multilinear forms $T_{i}$ of degree at most $d$. With a slight abuse of notation, the codimension of $V$ is the least number $m$ of multilinear forms needed. The density of $V$ is $|V|/|\mathbb{F}^{nd}|\in[0,1]$.
Theorem 3.4 ([Mil25]).
Any multilinear variety $V\subseteq(\mathbb{F}_{p}^{n})^{d}$ of density $\delta$ contains a multilinear variety of codimension at most $O_{d}(\log_{p}(1/\delta)+1)$.
We now give our short alternative proof to the Polynomial Gowers inverse conjecture in degree $d$, namely, the bound $\operatorname{cor}_{<d}(f)\geq\operatorname{bias}(\Delta^{d}f)^{O_{d}(1)}$ for every polynomial $f\in\operatorname{Poly}_{d}(\mathbb{F}_{p})$ with $d<p$.
Proof of $\operatorname{GI}_{d}(d)$ from Theorem 3.4.
Let $f\colon\mathbb{F}^{n}\to\mathbb{F}$ with $\deg(f)\leq d$ and $\operatorname{bias}(\Delta^{d}f)=\delta$, where $\mathbb{F}=\mathbb{F}_{p}$ and $d<p$. Recall that $\Delta^{d}f\colon(\mathbb{F}^{n})^{d}\to\mathbb{F}$ is a multilinear form of degree $d$, and let us write $\Delta^{d}f(x^{(1)},\ldots,x^{(d)})=\sum_{i=1}^{n}x^{(d)}_{i}T_{i}(x^{(1)},\ldots,x^{(d-1)})$. Put $\mathbf{T}=(T_{1},\ldots,T_{n})$, and consider the multilinear variety $V=\operatorname{Z}(\mathbf{T})\subseteq(\mathbb{F}^{n})^{d-1}$. Note that the density of $V$ is $\mathbb{P}(\mathbf{x}\in\operatorname{Z}(\mathbf{T}))=\operatorname{bias}(\Delta^{d}f)=\delta$. By Theorem 3.4, there are multilinear forms $\mathbf{S}=(S_{1},\ldots,S_{m})$, with each $S_{i}\colon(\mathbb{F}^{n})^{d-1}\to\mathbb{F}$ of degrees at most d1d-1, such that $\operatorname{Z}(\mathbf{S})\subseteq\operatorname{Z}(\mathbf{T})$ ($\subseteq\operatorname{Z}(\Delta^{d}f)$) and $m\leq O_{d}(\log_{p}(1/\delta)+1)$. We observe that for any nonzero $d$-linear form $T$, $\operatorname{AR}(T)\geq 2^{-d}$; this follows for example from Claim 3.3 in [Lov19]. Since we may assume $\Delta^{d}f\neq 0$, as otherwise $\delta=\operatorname{bias}(\Delta^{d}f)=1$ and there is nothing to prove, we have that $\log_{p}(1/\delta)=\operatorname{AR}(\Delta^{d}f)\geq 2^{-d}$. Thus, $\log_{p}(1/\delta)+1\leq(2^{d}+1)\log_{p}(1/\delta)$, and therefore
$m\leq O_{d}(\log_{p}(1/\delta)).$ (3)
Plugging in the diagonal, let $A_{i}(x)=S_{i}(x,\ldots,x)$, homogeneous with $\deg(A_{i})\leq\det(S_{i})\leq d-1$. Put $\mathbf{A}=(A_{1},\ldots,A_{m})$. We have $f_{d}(x)=(1/d!)\Delta^{d}f(x,\ldots,x)$ using $d<p$. Since $\operatorname{Z}(\mathbf{S})\subseteq\operatorname{Z}(\Delta^{d}f)$, we have that $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(f_{d})$. [^8] Indeed, $\operatorname{Z}(\mathbf{A})=\{x\,|\,\mathbf{A}(x)=0\}=\{x\,|\,\mathbf{S}(x,\ldots,x)=0\}\subseteq\{x\,|\,\Delta^{d}f(x,\ldots,x)=0\}=\operatorname{Z}(f_{d})$.
By Corollary 3.3, $\operatorname{cor}(f_{d},P)\geq p^{-dm}$ for some polynomial $P=\sum_{i=1}^{m}c_{i}A_{i}$ of degree at most d1d-1. Note that $\operatorname{cor}(f,\,P-f_{<d})=\operatorname{bias}(f-(P-f_{<d}))=\operatorname{bias}(f_{d}-P)$. Thus,
$\operatorname{cor}_{<d}(f)\geq\operatorname{bias}(f_{d}-P)\geq p^{-dm}\geq p^{-O_{d}(\log_{p}(1/\delta))}=\delta^{O_{d}(1)}$
using (3) in the last inequality. This completes the proof.
4 $\operatorname{rk}^{*}$
In this section we introduce our refined notion of rank, compare it to standard notions from the literature, and prove several properties that would be useful later.
Recall that for a form $f$, the rank of $f$ is the smallest number of reducible forms that sum to $f$. A natural extension of rank to general polynomials (that is, not necessarily homogeneous) expresses $f$ as a sum of reducible polynomials of degree at most $\deg(f)$ (so the factors are of degree below $\deg(f)$). Our new notion of rank does that, except it also adds a mild condition on the factors in the decomposition, or more specifically, on its factors of degree $1$. Intuitively, one can think of this as a very weak “quasirandomness” condition on the decomposition.
Call a set of degree-$1$ polynomials affinely independent if their linear components are linearly independent. For example, $x$ and $x+1$ are linearly independent but not affinely independent, since they have the same linear component, namely, $x$.
Definition 4.1 ($\operatorname{rk}^{*}$).
For a polynomial $f$, $\operatorname{rk}^{*}(f)$ is the smallest $r$ such that:
  • $f=\sum_{i=1}^{r}f_{i}$ for some reducible polynomials $f_{i}=\alpha_{i}\beta_{i}$ with $\deg(f_{i})\leq\deg(f)$, and
  • $\{\alpha_{i}\,|\,\deg(\alpha_{i})=1,\deg(f_{i})=\deg(f)\}$ are affinely independent, where $\deg(\alpha_{i})\leq\deg(\beta_{i})$.
We have $\operatorname{rk}^{*}(f)=\infty$ if no such decomposition exists, for example if $\deg(f)=1$; however, as $f$ need not be homogeneous, this can happen in higher degree as well (see Remark 4.2 below).
We note that if $f$ is homogeneous then $\operatorname{rk}^{*}(f)=\operatorname{rk}(f)$; indeed, the factors may be assumed to be linearly independent forms (see Lemma 4.6), so the affine independence condition is trivially satisfied.
As an example for $\operatorname{rk}^{*}$, consider a polynomial of the form $f=(A+B)\cdot\ell+B$, where $\deg(A)=\deg(B)>1$ and $\deg(\ell)=1$. While $f$ has the decomposition $f=\ell A+(\ell+1)B$, this is not a $\operatorname{rk}^{*}$-decomposition since its degree-$1$ factors $\ell$ and $\ell+1$ are not affinely independent. In fact, $\operatorname{rk}^{*}(f)$ may be arbitrarily large (see Remark 4.2 next).
Remark 4.2.
Interestingly, $\operatorname{rk}^{*}(f)$ can be more closely related to $\operatorname{bias}(f)$, or rather to analytic rank $\operatorname{AR}(f)=-\log_{p}(\operatorname{bias}(P))$, than other standard notions of rank in the literature. For concreteness, consider the notion used by Green and Tao in [GT09]: $\operatorname{rank}(f)$ is the least number $k$ of polynomials $\alpha_{1},\ldots,\alpha_{k}$ with $\deg(\alpha_{i})<\deg(f)$ such that $f\equiv\Gamma(\alpha_{1},\ldots,\alpha_{k})$ for some function $\Gamma\colon\mathbb{F}^{k}\to\mathbb{F}$. Consider the following polynomials $f$:
  • $f(x,y)=A(x)+\ell(y)$ with $\deg(A)>\deg(\ell)=1$. Clearly, $\operatorname{rank}(f)\leq\operatorname{rank}(A)+1<\infty$. However, $\operatorname{rk}^{*}(f)=\infty$, and moreover, $\operatorname{bias}(f)=\operatorname{bias}(A)\operatorname{bias}(\ell)=0$, so $\operatorname{AR}(f)=\infty$ as well.
  • $f(x,y)=(A-B)\ell+B$ with $\ell=\ell(y)$ of degree $1$, and $A=A(x)$,$B=B(x)$ random of degree $d\geq 2$ on $n$ variables. Clearly, $\operatorname{rank}(f)\leq 3$. In contrast, we claim that both $\operatorname{rk}^{*}(f)$ and $\operatorname{AR}(f)$ grow with $n$.
    Consider a $\operatorname{rk}^{*}$-decomposition of $f$. By affine independence, setting all degree-$1$ factors to 0 gives a non-empty affine subspace. Let $U$ be the non-empty intersection of this affine subspace with the hypersurface $\ell=c$, for an appropriate $c\in\mathbb{F}$. Then $\operatorname{codim}(U)\leq\operatorname{rk}^{*}(f)+1$, and the restriction $f^{\prime}$ of $f$ to $U$ has $\operatorname{rank}(f^{\prime})\leq\operatorname{rk}^{*}(f)$. By construction, $f^{\prime}$ is a nonzero linear combination of $A$ and $B$ restricted to $U$, so $\operatorname{rank}(f^{\prime})$ grows with $n$, and thus also $\operatorname{rk}^{*}(f)$.
    As for $\operatorname{AR}(f)$, we have $\operatorname{bias}(f)=\operatorname*{\mathbb{E}}_{x}\chi(B(x))\operatorname*{\mathbb{E}}_{y}\chi(\ell(y)(A(x)-B(x)))=\operatorname*{\mathbb{E}}_{x}\chi(B(x))\mathbf{1}_{A=B}(x)\\ =\sum_{b}\chi(b)\mathbb{P}(B=b,A=b)$, so $\operatorname{bias}(f)\approx(1/q)\sum_{b}\chi(b)\mathbb{P}(B=b)=(1/q)\operatorname{bias}(B)$, implying that $\operatorname{AR}(f)\approx n$, as claimed.
4.1 Properties of $\operatorname{rk}^{*}$
In general, $\operatorname{rk}^{*}$ is not subadditive, unlike other notions of rank. That is, the sum of two $\operatorname{rk}^{*}$-decompositions need not be another $\operatorname{rk}^{*}$-decomposition, for the simple reason that there could be linear dependencies between the linear forms in the degree-$1$ factors of the two decompositions. For example, the $\operatorname{rk}^{*}$ of $(x+1)R$ and of $xS$, where $\deg(R),\deg(S)\geq 2$, is $1$, but the $\operatorname{rk}^{*}$ of their sum $(x+1)R+xS$ need not be at most $2$; in fact, it can be unbounded (recall Remark 4.2).
That being said, $\operatorname{rk}^{*}$ is subadditive if the two polynomials are of different degrees.
Observation 4.3 (subadditivity).
$\operatorname{rk}^{*}(g+h)\leq\operatorname{rk}^{*}(g)+\operatorname{rk}^{*}(h)$ if $\deg(h)\neq\deg(g)$.
A basic fact about rank is that it is not inherited by partial derivatives: even if a form has low rank, a partial derivative might not. Consider, say, reducible forms; for example, we have $\operatorname{rk}(x\cdot S)\leq 1$ for any form $S=S(y)$, yet the partial derivative with respect to $x$ is $S(y)$, which may have arbitrarily high rank.
That being said, we show that, perhaps surprisingly, adding to a polynomial a partial derivative does not increase its $\operatorname{rk}^{*}$.
Lemma 4.4 (derivation invariance).
For any form $g$ and any direction vector $c$,
$\operatorname{rk}^{*}(g+\partial_{c}g)\leq 2\operatorname{rk}(g).$
Combining the above two properties, we show that $\operatorname{rk}^{*}$ satisfies a certain approximate “chain rule”, which would complement a chain rule for bias that we prove later in Section 6.
Proposition 4.5 (chain rule for $\operatorname{rk}^{*}$).
For any polynomial $g+h$ with top-degree component $g$, and any direction vector $c$,
$\operatorname{rk}^{*}(g+h)\leq 2\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g).$
We now prove the above properties of $\operatorname{rk}^{*}$.
Proof of Observation 4.3.
Put $r_{1}=\operatorname{rk}^{*}(g)$, $r_{2}=\operatorname{rk}^{*}(h)$. Let $g=\sum_{i=1}^{r_{1}}g_{i}$ and $h=\sum_{i=1}^{r_{2}}h_{i}$, with $g_{i}$ and $h_{i}$ reducible polynomials with $\deg(g_{i})\leq\deg(g)$ and $\deg(h_{i})\leq\deg(h)$, be decompositions witnessing $\operatorname{rk}^{*}(g)$ and $\operatorname{rk}^{*}(h)$, respectively. We have $g+h=\sum_{i=1}^{r_{1}}g_{i}+\sum_{i=1}^{r_{2}}h_{i}$. Assume without loss of generality that $\deg(h)<\deg(g)$. Since $\deg(h_{i})\leq\deg(h)$, observe that the affine independence condition for $\operatorname{rk}^{*}(g+h)$ applies only for the $g_{i}$, and is therefore satisfied by the definition of $\operatorname{rk}^{*}(g)$. Thus, $\operatorname{rk}^{*}(g+h)\leq r_{1}+r_{2}=\operatorname{rk}^{*}(g)+\operatorname{rk}^{*}(h)$, as needed. ∎
Lemma 4.6.
For every $g\in\operatorname{Form}$ and a shortest $\operatorname{rk}$-decomposition $g=\sum_{i}\alpha_{i}\beta_{i}$, the forms $(\alpha_{i})_{i}$ (and $(\beta_{i})_{i}$) are linearly independent.
Proof.
Put $r=\operatorname{rk}(g)$, and let $g=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ be a $\operatorname{rk}$-decomposition. Suppose for contradiction that $\alpha_{r}=\sum_{i=1}^{r-1}c_{i}\alpha_{i}$ with $c_{i}$ scalars. By homogeneity, we may assume that $\deg(\alpha_{i})=\deg(\alpha_{r})$ whenever $c_{i}\neq 0$. Observe that $g$ has the decomposition $g=\sum_{i=1}^{r-1}\alpha_{i}(\beta_{i}+c_{i}\beta_{r})$. Each summand is a reducible form, since $\alpha_{i}$ is homogeneous as well as $\beta_{i}+c_{i}\beta_{r}$; indeed, $\deg(\beta_{i})=k-\deg(\alpha_{i})=k-\deg(\alpha_{r})=\deg(\beta_{r})$ whenever $c_{i}\neq 0$. Thus, $r=\operatorname{rk}(g)\leq r-1$, a contradiction. ∎
Next we prove the invariance of $\operatorname{rk}^{*}$ with respect to adding a derivative.
Proof of Lemma 4.4.
Put $r=\operatorname{rk}(g)$ and $k=\deg(g)$, and write $g=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ with forms $\alpha_{i},\beta_{i}$ satisfying $\deg(\alpha_{i})\leq\deg(\beta_{i})<k$ and $\deg(\alpha_{i}\beta_{i})=k$.
By the product rule, $\partial g_{c}=\sum_{i=1}^{r}\alpha_{i}(\partial_{c}\beta_{i})+(\partial_{c}\alpha_{i})\beta_{i}$. Observe that we have the following fortunate identity:
$g+\partial_{c}g=\sum_{i=1}^{r}(\alpha_{i}+\partial_{c}\alpha_{i})(\beta_{i}+\partial_{c}\beta_{i})-\sum_{i=1}^{r}(\partial_{c}\alpha_{i})(\partial_{c}\beta_{i}).$
Now, since $\deg(\alpha_{i}+\partial_{c}\alpha_{i})=\deg(\alpha_{i})\leq\deg(\beta_{i})=\deg(\beta_{i}+\partial_{c}\beta_{i})$, the linear components of the factors $\alpha_{i}+\partial_{c}\alpha_{i}$ are precisely the linear factors $\alpha_{i}$ in the above decomposition for $g$, which by Lemma 4.6 are linearly independent. As for the remaining terms, note that $\deg((\partial_{c}\alpha_{i})(\partial_{c}\beta_{i}))\leq\deg(\alpha_{i}\beta_{i})-2=k-2<\deg(g+\partial_{c}g)$, so $\operatorname{rk}^{*}(g+\partial_{c}g)$ has no requirement on their (linear) factors. We deduce that the above (non-homogeneous) decomposition is a $\operatorname{rk}^{*}$-decomposition, so $\operatorname{rk}^{*}(g+\partial_{c}g)\leq 2r=2\operatorname{rk}(g)$, as needed.
We are now ready to prove our chain rule for $\operatorname{rk}^{*}$.
Proof of Proposition 4.5.
We have
$\operatorname{rk}^{*}(g+h)$ $=\operatorname{rk}^{*}\big((g+\partial_{c}g)+(h-\partial_{c}g)\big)$
$\leq\operatorname{rk}^{*}(g+\partial_{c}g)+\operatorname{rk}^{*}(h-\partial_{c}g)\leq 2\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g)$
where the first inequality uses the subadditivity of $\operatorname{rk}^{*}$ in Observation 4.3, and the second inequality uses Lemma 4.4. ∎
4.2 Lower-degree correlation
Recall that standard correlation results show that if a polynomial $f$ has a short decomposition in terms of lower degree polynomials, then it has a significant correlation with a polynomial of degree at most $\deg(f)-1$. In this section we show that if a polynomial $f$ has a short $\operatorname{rk}^{*}$-decomposition, then it has a significant correlation with a polynomial of degree at most $\deg(f)-2$. This relies on our correlation lemma in Lemma 3.1, but here we only assume small $\operatorname{rk}^{*}$.
Proposition 4.7 (lower-degree correlation).
For any $f\in\operatorname{Poly}_{k}(\mathbb{F}_{q})$ with $k\geq 3$,
$\operatorname{cor}_{<k-1}(f)\geq q^{-2\operatorname{rk}^{*}(f)}.$
In order to convert a short $\operatorname{rk}^{*}$-decomposition to significant correlation, via Lemma 3.1, we use the following lemma which gives an essentially optimal lower bound on the joint zero probability $\mathbb{P}(\operatorname{Z}(\mathbf{A}))$.
Lemma 4.8 (perturbation lemma).
For any $f\in\operatorname{Poly}_{k}(\mathbb{F}_{q})$ with $k\geq 3$, we have $f-c\in\langle\mathbf{A}\rangle$, where $\mathbf{A}=(A_{1},\ldots,A_{m})$ and $c\in\mathbb{F}_{q}$, such that:
  • $1\leq\deg(A_{i})\leq k-2$,
  • $\mathbb{P}(\operatorname{Z}(\mathbf{A}))\geq q^{-m}$,
  • and $m\leq 2\operatorname{rk}^{*}(f)$.
Proof.
Put $r=\operatorname{rk}^{*}(P)$ and $\mathbb{F}=\mathbb{F}_{q}$. Write $f=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ with $1\leq\deg(\alpha_{i})\leq\deg(\beta_{i})\leq k-1$ and $\deg(\alpha_{i}\beta_{i})\leq k$. Assume without loss of generality that all polynomials $\alpha_{i}$ satisfying both $\deg(\alpha_{i})=1$ and $\deg(\alpha_{i}\beta_{i})=k$ are precisely $\alpha_{t+1},\ldots,\alpha_{r}$ for some $t\leq r$. Observe that for every $i\leq t$ we necessarily have $\deg(\beta_{i})=\deg(\alpha_{i}\beta_{i})-\deg(\alpha_{i})\leq k-2$.
Put $\mathbf{A}^{1}=(\alpha_{1},\ldots,\alpha_{t})$, $\mathbf{A}^{2}=(\alpha_{t+1},\ldots,\alpha_{r})$, and $\mathbf{B}^{1}=(\beta_{1},\ldots,\beta_{t})$. Since
$\sum_{y,z\in\mathbb{F}^{t}}\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y,z)\,|\,\mathbf{A}^{2}=0\big)=1,$
there exist $y^{1},z^{1}\in\mathbb{F}^{t}$ such that
$\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y^{1},z^{1})\,|\,\mathbf{A}^{2}=0\big)\geq q^{-2t}.$ (4)
Put $y=(y^{1},y^{2})\in\mathbb{F}^{r}$ with $y^{2}=0\in\mathbb{F}^{r-t}$. We decompose $f$ as follows (recall $y_{i}=0$ for $i\geq t$):
$f=\sum_{i=1}^{r}(\alpha_{i}-y_{i})\beta_{i}+\sum_{i=1}^{t}y_{i}(\beta_{i}-z_{i})+\sum_{i=1}^{t}y_{i}z_{i}.$
Let $A_{i}=\alpha_{i}-y_{i}$ for $i\leq r$, let $A_{r+i}=\beta_{i}-z_{i}$ for $i\leq t$, and let $c=\sum_{i=1}^{t}y_{i}z_{i}$ ($\in\mathbb{F}$). Then $f-c\in\langle A_{1},...,A_{m}\rangle$ with $m=r+t\leq 2r$. Moreover, for $i\leq k$ we have $\deg(A_{r+i})=\deg(\beta_{i})\leq k-2$, and for $i\leq r$ we have $\deg(A_{i})=\deg(\alpha_{i})\leq\lfloor k/2\rfloor\leq k-2$, using our assumption $k\geq 3$.
Finally, by the definition of $\operatorname{rk}^{*}(f)$, the linear components of $\alpha_{t+1},\ldots,\alpha_{r}$ are linearly independent. Thus, the affine map $\mathbf{A}^{2}\colon\mathbb{F}^{n}\to\mathbb{F}^{r-t}$ is perfectly equidistributed, and in particular, $\mathbb{P}(\mathbf{A}^{2}=0)\geq q^{-(r-t)}$. Therefore,
$\mathbb{P}(\forall i\colon A_{i}=0)$ $=\mathbb{P}((\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})=(y^{1},y^{2},z^{1}))$
$=\mathbb{P}(\mathbf{A}^{2}=0)\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y^{1},z^{1})\,|\,\mathbf{A}^{2}=0\big)\geq q^{-(r-t)}q^{-2t}=q^{-m},$
where the inequality combines the equidistribution of $\mathbf{A}^{2}$ together with (4). This completes the proof. ∎
Proof of Proposition 4.7.
Apply Lemma 4.8 on $f$ to obtain $m$ polynomials $\mathbf{A}=(A_{1},...,A_{m})\in\operatorname{Poly}_{k-2}^{m}(\mathbb{F}_{q})$ and a scalar $c_{0}\in\mathbb{F}_{q}$ such that $f-c_{0}\in\langle\mathbf{A}\rangle$, where $\mathbb{P}(\operatorname{Z}(\mathbf{A}))\geq q^{-m}$ and $m\leq 2\operatorname{rk}^{*}(f)$. Now apply Lemma 3.1 on $f,A_{1},\ldots,A_{m}$ to deduce that
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\Big|\operatorname{\mathfrak{bias}}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)\Big|$ $\geq\Big|\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)\Big|$
[…]
Since $\sum_{i=1}^{m}c_{i}A_{i}\in\operatorname{Poly}_{k-2}$, we deduce that
$\operatorname{cor}_{<k-1}(f)\geq\max_{c\in\mathbb{F}^{m}}\operatorname{cor}\Big(f,\,\sum_{i=1}^{m}c_{i}A_{i}\Big)\geq q^{-2\operatorname{rk}^{*}(f)}.\qed$
5 Discrete derivatives
For a multilinear form $T$, we denote $\partial_{c}T(\mathbf{v})=T(\mathbf{v},c)$. In other words, $\partial_{c}T$ is the directional derivative with respect to the (say) last vector variable. We denote $\boldsymbol{\nabla}T(\mathbf{v})=\big(T(\mathbf{v},e_{i})\big)_{i}$, so that $\boldsymbol{\nabla}T$ is the gradient with respect to the last vector variable and $\partial_{c}T=c\cdot\boldsymbol{\nabla}T$.
Our main result in this section is a particularly useful description of the order-$d$ discrete derivative of forms of degree $d+1$.
Proposition 5.1.
For any $g\in\operatorname{Form}_{d+1}(\mathbb{F})$, where $\operatorname{ch}(\mathbb{F})\neq 2$, and any $\mathbf{v}\in(\mathbb{F}^{n})^{d}$,
$\Delta^{d}_{\mathbf{v}}g(x)=\overline{g}(\mathbf{v},x^{\prime})\quad\text{ where }x^{\prime}=x+\frac{1}{2}\sum_{t=1}^{d}v^{(t)}.$
Put differently, $\Delta^{d}_{\mathbf{v}}g(x)=x^{\prime}\cdot\boldsymbol{\nabla}\overline{g}(\mathbf{v})$. We will also show that, although not entirely obvious, polarization and formal derivation commute.
Lemma 5.2.
For any form $g$ and any direction vector $c$,
$\partial_{c}\overline{g}=\overline{\partial_{c}g}.$
It is instructive to consider simple examples of these somewhat subtle identities.
Example 5.3 (Example for Lemma 5.2).
For $g(x)=x_{1}x_{2}^{2}$, we have $\partial_{c}g(x)=c\cdot(x_{2}^{2},\,2x_{1}x_{2})$, so that $\overline{\partial_{c}g}(x,y)=c\cdot(2x_{2}y_{2},\,2x_{1}y_{2}+2x_{2}y_{1})$. On the other hand, $\overline{g}(x,y,z)=2x_{1}y_{2}z_{2}+2y_{1}x_{2}z_{2}+2z_{1}x_{2}y_{2}$, so that $\partial_{c}\overline{g}(x,y)=\overline{g}(x,y,c)=c\cdot(2x_{2}y_{2},\,2x_{1}y_{2}+2x_{2}y_{1})$ as well.
Example 5.4 (Example for Proposition 5.1).
For $d=2$ and $g(x)=x_{1}x_{2}x_{3}$, we have by definition
$\Delta^{2}_{u,v}g(x)$ $=g(x+u+v)-g(x+u)-g(x+v)+g(x)$
$=(x_{1}+u_{1}+v_{1})(x_{2}+u_{2}+v_{2})(x_{3}+u_{3}+v_{3})$
$-(x_{1}+u_{1})(x_{2}+u_{2})(x_{3}+u_{3})-(x_{1}+v_{1})(x_{2}+v_{2})(x_{3}+v_{3})+x_{1}x_{2}x_{3}.$
One can group the terms above in the following suggestive way:
$\Delta^{2}_{u,v}g(x)$ $=x_{1}(u_{2}v_{3}+v_{2}u_{3})+x_{2}(u_{1}v_{3}+v_{1}u_{3})+x_{3}(u_{1}v_{2}+v_{1}u_{2})$
$+(u_{1}u_{2}v_{3}+u_{1}v_{2}u_{3}+v_{1}u_{2}u_{3})+(v_{1}v_{2}u_{3}+v_{1}u_{2}v_{3}+u_{1}v_{2}v_{3}),$
and a short inspection reveals
$\Delta^{2}_{u,v}g(x)$ $=\big(x+\frac{1}{2}u+\frac{1}{2}v\big)\cdot\big(u_{2}v_{3}+v_{2}u_{3},\,u_{1}v_{3}+v_{1}u_{3},\,u_{1}v_{2}+v_{1}u_{2}\big).$
Note that by Lemma 5.2, we can compute $\boldsymbol{\nabla}\overline{g}$ as $\overline{\boldsymbol{\nabla}g}$. And indeed,
$\Delta^{2}_{u,v}g(x)=\big(x+\frac{1}{2}u+\frac{1}{2}v\big)\cdot\overline{\big(x_{2}x_{3},\,x_{1}x_{3},\,x_{1}x_{2}\big)}(u,v)\;.$
5.1 Discrete derivatives: Proofs
We first give a combinatorial description of the iterated discrete derivatives of any form (and by linearity, of any polynomial) in terms of semi-surjections. Henceforth,
$\varphi\colon[k]\twoheadrightarrow[d]\text{ means }\varphi\colon[k]\to[d+1]\text{ with }\operatorname{Im}(\varphi)\supseteq[d].$
We refer to a function $\varphi\colon[k]\twoheadrightarrow[d]$ as a semi-surjection since it is surjective on its codomain except, perhaps, the last element, $d+1$. Denote by $S_{n}$ the set of permutations $\sigma\colon[n]\to[n]$.
Lemma 5.5.
For any form $f(x)=\sum_{I}c_{I}\prod_{j=1}^{k}x_{I_{j}}$, with each $I=(I_{1},\ldots,I_{k})$ a sorted $k$-tuple of variables and $c_{I}$ a scalar,
$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})=\sum_{I}c_{I}\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{I_{j}}^{(\varphi(j))}.$
Proof.
By linearity we may assume that $f$ is a single monomial, $f=\prod_{j=1}^{k}x_{i_{j}}$, so it suffices to prove
$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})=\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{i_{j}}^{(\varphi(j))}.$
We proceed by induction on $d\geq 0$. The base case $d=0$ holds since the only function $\varphi\colon[k]\to\{1\}$ is $\varphi=1$, so the right hand side is $\prod_{j=1}^{k}x_{i_{j}}^{(1)}=f(x^{(1)})=\Delta^{0}_{\mathbf{x}}f(x^{(1)})$.
For the induction step, we have
$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$ $=\Delta_{x^{(d)}}\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)})=\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)}+x^{(d)})-\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)})$
$=\sum_{\varphi\colon[k]\twoheadrightarrow[d-1]}\,\prod_{j\colon\varphi(j)\neq d}x_{i_{j}}^{(\varphi(j))}\Big(\prod_{j\colon\varphi(j)=d}(x_{i_{j}}^{(d+1)}+x_{i_{j}}^{(d)})-\prod_{j\colon\varphi(j)=d}x_{i_{j}}^{(d+1)}\Big)$
$=\sum_{\varphi\colon[k]\twoheadrightarrow[d-1]}\,\prod_{j\colon\varphi(j)\neq d}x_{i_{j}}^{(\varphi(j))}\sum_{\begin{subarray}{c}\tau\colon\varphi^{\raisebox{0.43054pt}{$\scriptscriptstyle\!-\!1$}}(d)\to\{d,d+1\}\colon\\ d\in\operatorname{Im}(\tau)\end{subarray}}\,\prod_{j\colon\varphi(j)=d}x_{i_{j}}^{(\tau(j))}$
$=\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{i_{j}}^{(\varphi(j))},$
where the third equality uses the induction hypothesis, and the last equality uses the natural bijection between semi-surjections $\varphi^{\prime}\colon[k]\twoheadrightarrow[d]$ and pairs of semi-surjections $(\varphi,\tau)$ where $\varphi\colon[k]\twoheadrightarrow[d-1]$ and $\tau\colon\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle\!-\!1$}}(d)\to\{d,d+1\}$ with $d\in\operatorname{Im}(\tau)$. This completes the induction step and the proof.
We record two facts about polarizations, which can be deduced from Lemma 5.5 effortlessly.
Fact 5.6.
For any monomial $g(x)=\prod_{j=1}^{k}x_{i_{j}}$,
$\overline{g}(x^{(1)},\ldots,x^{(k)})=\sum_{\pi\in S_{k}}\,\prod_{j=1}^{k}x_{i_{j}}^{(\pi(j))}\,.$
Fact 5.7.
Any $g\in\operatorname{Form}_{k}$ satisfies $\overline{g}(x,x,\ldots,x)=k!g(x)$.
Using the explicit description of polarization, we now prove that polarization commutes with the formal derivative.
Lemma (5.2).
For any form $g$ and direction vector $c$,
$\partial_{c}\overline{g}=\overline{\partial_{c}g}.$
Proof of Lemma 5.2.
It suffices to show $\partial_{x_{n}^{(k)}}\overline{g}(\mathbf{x})=\overline{\partial_{x_{n}}g}(\mathbf{x})$ for $g(x)=\prod_{j=1}^{k}x_{i_{j}}$ ($i_{1}\leq\cdots\leq i_{k}$) a single monomial, and $\mathbf{x}=(x^{(1)},\ldots,x^{(k)})$, since the result would follow by symmetry and linearity. On the one hand, [^9] For example, $\partial_{x_{2}}(x_{1}x_{2}^{2})=x_{1}x_{2}+x_{1}x_{2}$.
$\partial_{x_{n}}g(x)=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\prod_{\begin{subarray}{c}j=1\\ j\neq t\end{subarray}}^{k}x_{i_{j}}\,=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\prod_{j=1}^{k-1}x_{i_{j}},$
and therefore
$\overline{\partial_{x_{n}}g}(x^{(1)},\ldots,x^{(k-1)})=\sum_{\sigma\in S_{k-1}}\,\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\prod_{j=1}^{k-1}x_{i_{j}}^{(\sigma(j))}.$ (5)
On the other hand,
$\overline{g}(x^{(1)},\ldots,x^{(k)})=\sum_{\sigma\in S_{k}}\,\prod_{j=1}^{k}x_{i_{j}}^{(\sigma(j))}$
is multilinear, so its only monomials containing $x^{(k)}_{n}$ are those where an occurrence of $x_{n}$ is replaced by $x_{n}^{(k)}$; therefore,
$\partial_{x^{(k)}_{n}}\overline{g}(x^{(1)},\ldots,x^{(k)})=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\sum_{\begin{subarray}{c}\sigma\in S_{k}\\ \pi(t)=k\end{subarray}}\,\prod_{\begin{subarray}{c}j=1\\ j\neq t\end{subarray}}^{k}x_{i_{j}}^{(\sigma(j))}=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\sum_{\begin{subarray}{c}\sigma\in S_{k}\\ \pi(t)=k\end{subarray}}\,\prod_{j=1}^{k-1}x_{i_{j}}^{(\sigma(j))}\;.$ (6)
A quick inspection shows that (5) and (6) are equal, as needed.
We are now ready to prove our formula for the order-$d$ derivative of forms of degree $d+1$.
Proof of Proposition 5.1.
By linearity, we may assume that $f$ is a single monomial, $f=\prod_{j=1}^{d+1}x_{i_{j}}$, of degree $d+1$. For a function $\varphi\colon[k]\to[d+1]$, we abbreviate $f^{(\varphi)}=\prod_{j=1}^{d+1}x_{i_{j}}^{(\varphi(j))}$. By Lemma 5.5,
$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$ $=\sum_{\varphi\colon[d+1]\twoheadrightarrow[d]}\,f^{(\varphi)}$
$=\sum_{\sigma\in S_{d+1}}\,f^{(\sigma)}\,+\,\sum_{t=1}^{d}\,\sum_{\varphi\in S_{d}^{t}}\,f^{(\varphi)}$
where $S_{d}^{t}$ denotes the set of all surjections $\varphi\colon[d+1]\to[d]$ satisfying $|\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle\!-\!1$}}(t)|=2$. We claim that for any $t\in[d]$, there is a 22-to-11 map from $S_{d+1}$ to $S_{d}^{t}$. Indeed, we map each permutation $\sigma\in S_{d+1}$ to the function $\sigma_{t}\in S_{d}^{t}$ given by
$\sigma_{t}(x):=\begin{cases}t&\text{if }\sigma(x)=d+1\\ \sigma(x)&\text{else}\end{cases}\;.$
Put differently, $\sigma_{t}(x)=\sigma_{\sigma^{\raisebox{0.43054pt}{$\scriptscriptstyle\!-\!1$}}(d+1)\to t}$, where for any function $h$ we write $h_{x_{0}\to y_{0}}$ for
$h_{x_{0}\to y_{0}}(x)=\begin{cases}y_{0}&\text{if }x=x_{0}\\ h(x)&\text{else}\end{cases}\;.$
To see that this map is 22-to-11, let $\varphi\in S_{d}^{t}$, and consider the two functions $\varphi_{a\mapsto d+1}$ and $\varphi_{b\mapsto d+1}$. That both functions are in $S_{d+1}$ follows from our assumption on $\varphi$. A moment’s thought reveals that both—and none others—are mapped to $\varphi$, that is, $\sigma_{\sigma^{\raisebox{0.43054pt}{$\scriptscriptstyle\!-\!1$}}(d+1)\to t}=\varphi$ if and only if $\sigma=\varphi_{a\mapsto d+1}$ or $\sigma=\varphi_{b\mapsto d+1}$, proving our claim. Going back to $\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$, we thus have
$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$ $=\sum_{\sigma\in S_{d+1}}\,f^{(\sigma)}\,+\,\sum_{t=1}^{d}\,\frac{1}{2}\sum_{\sigma\in S_{d+1}}f^{(\sigma_{t})}$
$=\overline{f}(\mathbf{x},x^{(d+1)})+\sum_{t=1}^{d}\,\frac{1}{2}\overline{f}(\mathbf{x},x^{(t)})$
$=\overline{f}\Big(\mathbf{x},x^{(d+1)}+\frac{1}{2}\sum_{t=1}^{d}x^{(t)}\Big).$
using the multilinearity of $\overline{f}$ in the last equality. This completes the proof. ∎
6 Proof of Theorem 1
6.1 Bias of general polynomials
While the bias of a multilinear form is roughly characterized by its rank (recall Theorem 2.1), the situation is more involved for general, non-homogeneous polynomials.
Remark 6.1.
While the bias of a polynomial is closely related to the bias of its top-degree component, it generally has little to do with the bias of its lower-degree components. For example, consider the univariate polynomial $x^{2}+x$, over $\mathbb{F}_{p}$ with $p>2$. We have
$\operatorname{bias}(x^{2}+x)\cdot p=\Big|\sum_{x\in\mathbb{F}_{p}}\omega_{p}^{x^{2}+x}\Big|=\Big|\omega_{p}^{-1/4}\sum_{x\in\mathbb{F}_{p}}\omega_{p}^{(x+1/2)^{2}}\Big|=\Big|\sum_{y\in\mathbb{F}_{p}}\omega_{p}^{y^{2}}\Big|=\sqrt{p},$
where the last step uses the formula for a basic quadratic Gauss sum. Thus, $\operatorname{bias}(x^{2}+x)=\operatorname{bias}(x^{2})=p^{-1/2}$, whereas for the lower degree component of $x^{2}+x$ we have $\operatorname{bias}(x)=0$.
Remark 6.2.
[…]
First, we prove a rather general chain rule for bias. For an event $E$, we henceforth define the conditional bias as $\operatorname{\mathfrak{bias}}(f\,|\,E)=\operatorname*{\mathbb{E}}(\chi(f)\,|\,E)$.
Lemma 6.3 (Chain rule for bias, a general form).
Let $A(x,y)$ and $B(y)$ be functions over a finite field, and assume $A$ is linear in $x$. For the event $E=\{y|\forall x\colon A(x,y)=0\}$, we have
$\operatorname{\mathfrak{bias}}(A+B)=\operatorname{\mathfrak{bias}}(A)\operatorname{\mathfrak{bias}}(B\,|\,E).$
Proof.
Observe that for any fixed $y\in E^{c}$, $A(x,y)+B(y)$ is polynomial in $x$ of degree exactly $1$, and thus takes any value in the field with the same probability. Furthermore observe that
$\operatorname{bias}(A)=\operatorname*{\mathbb{E}}(\chi(A))=\operatorname*{\mathbb{E}}_{y}\operatorname*{\mathbb{E}}_{x}(\chi(A(x,y)))=\operatorname*{\mathbb{E}}_{y}\mathbf{1}_{E}(y)=\mathbb{P}(E).$
Therefore, by the law of total expectation,
$\operatorname{bias}(A+B)$ $=\operatorname*{\mathbb{E}}(\chi(A+B))=\mathbb{P}(E)\operatorname*{\mathbb{E}}(\chi(A+B)\,|\,E)+\mathbb{P}(E^{c})\operatorname*{\mathbb{E}}(\chi(A+B)\,|\,E^{c})$
$=\operatorname{bias}(A)\operatorname{bias}(B\,|\,E)+\mathbb{P}(E^{c})\cdot 0.\qed$
Our main result in this subsection bounds the bias of the $d$-times iterated derivative of $f\in\operatorname{Poly}_{d+1}$ in terms of the components of $f$.
Proposition 6.4.
For $f\in\operatorname{Poly}_{d+1}(\mathbb{F})$ where $\operatorname{ch}(\mathbb{F})\neq 2$, with $g=f_{d+1}$ and $h=f_{d}$,
$\operatorname{bias}(\Delta^{d}f)\leq\operatorname*{\mathbb{E}}_{c}\operatorname{bias}(\overline{h}-\partial_{c}\overline{g})\quad\text{ and }\quad\operatorname{bias}(\Delta^{d}f)\leq\operatorname{bias}(\overline{g}).$
Proof.
By Proposition 5.1, $\Delta^{d}_{\mathbf{v}}g(x)=\overline{g}(\mathbf{v},x^{\prime})$ where $x^{\prime}=x+\frac{1}{2}\sum_{t=1}^{d}v^{(t)}$. Thus, $\Delta^{d}_{\mathbf{v}}f(x)=\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v})$, and therefore
$\operatorname{\mathfrak{bias}}(\Delta^{d}_{\mathbf{v}}f(x))=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v}))=\operatorname*{\mathbb{E}}_{x^{\prime}}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v}))=\operatorname{\mathfrak{bias}}(\overline{g}+\overline{h}).$
This has two implications. First, it follows from Lemma 6.3 that
$\operatorname{bias}(\overline{g}+\overline{h})\leq\operatorname{bias}(\overline{g}).$
Second, we have
$\operatorname{\mathfrak{bias}}(\overline{g}+\overline{h})$ $=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x)+\overline{h}(\mathbf{v}))=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\partial_{x}\overline{g}(\mathbf{v})+\overline{h}(\mathbf{v}))$
$=\operatorname*{\mathbb{E}}_{c}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\partial_{-c}\,\overline{g}(\mathbf{v})+\overline{h}(\mathbf{v}))=\operatorname*{\mathbb{E}}_{c}\operatorname{\mathfrak{bias}}(\overline{h}-\partial_{c}\overline{g}).$
These complete the proof. ∎
6.2 Main proof
We now put everything together and prove Theorem 1.
Theorem (1).
For every $f\in\operatorname{Poly}_{d+1}(\mathbb{F}_{q})$, with $\operatorname{ch}(\mathbb{F})>d+1$,
$\operatorname{cor}_{<d}(f)\geq\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}.$
Proof.
For $d=1$ there is nothing to prove, since $\operatorname{cor}_{<1}(f)^{2}=\operatorname{bias}(\Delta f)$; indeed, $\operatorname{bias}(f)^{2}=\operatorname{bias}(\Delta f)\geq\operatorname{cor}_{<1}(f)^{2}\geq\operatorname{cor}(f,0)^{2}=\operatorname{bias}(f)^{2}$ using (1). We thus assume $d\geq 2$. [^11] Recall that polynomial bounds for the Gowers inverse theorem are known also for $d=2$ (via Fourier analysis; e.g., Proposition 2.2 in [GT08]) and $d=3$ [GG+24, GG+25]. However, our proof does not rely on any of these.
Let $g=f_{d+1}$ and $h=f_{d}$ be the components of $f$ of degrees $d+1$ and $d$, respectively. We first claim that, for any direction vector $c$,
$\operatorname{rk}(g)\leq\operatorname{rk}(\overline{g})\quad\text{ and }\quad\operatorname{rk}(h-\partial_{c}g)\leq\operatorname{rk}(\overline{h}-\partial_{c}\overline{g}).$ (7)
To see this, abbreviate $\mathbf{x}^{t}=(x,\ldots,x)$ ($t$ times), and observe that $\partial_{c}\overline{g}(\mathbf{x}^{d})=\overline{\partial_{c}g}(\mathbf{x}^{d})=d!\partial_{c}g(x)$, using Lemma 5.2 and Fact 5.7. Moreover by Fact 5.7, we have $\overline{h}(\mathbf{x}^{d})=d!h(x)$ and $\overline{g}(\mathbf{x}^{d+1})=(d+1)!g(x)$. As $\operatorname{ch}(\mathbb{F})>d+1$, we thus have $g(x)=\overline{g}(\mathbf{x}^{d})/(d+1)!$ and $(h-\partial_{c}g)(x)=(\overline{h}-\partial_{c}\overline{g})(\mathbf{x}^{d})/d!$, proving (7).
By Proposition 6.4, $\operatorname{bias}(\Delta^{d}f)\leq\operatorname{bias}(\overline{g})$, and $\operatorname{bias}(\Delta^{d}f)\leq\operatorname{bias}(\overline{h}-\partial_{c}\overline{g})$ for some direction vector $c$ which we henceforth fix. Put $\operatorname{bias}(\Delta^{d}f)=q^{-r}$, so that
$\operatorname{bias}(\overline{g})\geq q^{-r}\quad\text{ and }\quad\operatorname{bias}(\overline{h}-\partial_{c}\overline{g})\geq q^{-r}.$ (8)
Now, combine (7), (8), and Theorem 2.1 to deduce that
$\operatorname{rk}(g)\leq\operatorname{rk}(\overline{g})\leq\tilde{O}_{d}(r)\quad\text{ and }\quad\operatorname{rk}(h-\partial_{c}g)\leq\operatorname{rk}(\overline{h}-\partial_{c}\overline{g})\leq\tilde{O}_{d}(r).$
Apply Proposition 4.5 to obtain
$\operatorname{rk}^{*}(g+h)\leq 2\operatorname{rk}(g)+\operatorname{rk}(h-\partial_{c}g)\leq\tilde{O}_{d}(r).$
Next, apply Proposition 4.7 with $k=d+1$ ($\geq 3$ by assumption) on $g+h\in\operatorname{Poly}_{d+1}$. We deduce that there exists a polynomial $P$ with $\deg(P)\leq k-2=d-1$ such that
$\operatorname{cor}(g+h,\,P)\geq q^{-2\operatorname{rk}^{*}(g+h)}\geq 1/q^{\tilde{O}_{d}(r)}.$
To finish, write $f=g+h+f_{<d-1}$. We claim that $P+f_{<d-1}\in\operatorname{Poly}_{d-1}$ significantly correlates with $f$; indeed,
$\operatorname{cor}_{<d}(f)$ $\geq\operatorname{cor}(f,\,P+f_{<d-1})=\operatorname{bias}(f-(P+f_{<d-1}))=\operatorname{bias}(g+h-P)$
$=\operatorname{cor}(g+h,\,P)\geq 1/q^{\tilde{O}_{d}(r)}=\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}.$
This completes the proof. ∎
Acknowledgment.
We thank Daniel G. Zhu for insightful comments on an earlier draft.

### 30 Apr 2026 v2

@@
 We moreover provide a nearly polynomial inverse theorem for homogeneous polynomials of any degree smaller than $2d$.
+Our methods may be of independent interest, and include a refined notion of polynomial decomposition that captures correlation with polynomials of lower degree than classical notions do, and a new correlation lemma that improves upon similar lemmas in the literature.
+Additionally, we illustrate the usefulness of the new correlation lemma by using it to give an alternative proof for the aforementioned result of Milićević and Rand¯elović.
 1 Introduction
-Let $\mathbb{F}=\mathbb{F}_{p}$ be any prime finite field.
+Let $\mathbb{F}$ be a finite field, for simplicity prime $\mathbb{F}=\mathbb{F}_{p}$.
 The discrete derivative of a function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ in direction $v\in\mathbb{F}^{n}$ at the point $x\in\mathbb{F}^{n}$ is $\Delta_{v}f(x)=f(x+v)-f(x)$.
-More generally, the (discrete) derivative of order $d$, in directions $\mathbf{v}=(v^{(1)},\ldots,v^{(d)})\in(\mathbb{F}^{n})^{d}$ at the point $x\in\mathbb{F}^{n}$, is
+More generally, the (discrete) derivative of order $d$, in directions $\mathbf{v}=(v^{(1)},\ldots,v^{(d)})\in(\mathbb{F}^{n})^{d}$ at $x\in\mathbb{F}^{n}$, is
 $\Delta^{d}_{\mathbf{v}}f(x):=\Delta_{v^{(d)}}\cdots\Delta_{v^{(1)}}f(x).$
-A function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ is a polynomial of degree below $d$, denoted $f\in\operatorname{Poly}_{<d}$, if and only if $\Delta^{d}f=0$ (that is, $\Delta^{d}_{\mathbf{v}}f(x)=0$ for every $(x,\mathbf{v})\in(\mathbb{F}^{n})^{d+1}$).
-An important (qualitative) result about polynomials over $\mathbb{F}$, of characteristic $\operatorname{ch}(\mathbb{F})>d$, is that the derivative characterization is robust: $f$ is correlated with $\operatorname{Poly}_{<d}$ if and only if $\Delta^{d}f$ is biased.
-This lies at the core of the theory of Gowers norms [Gow01, Tao12, HHL19], and is usually formulated using exponential sums: [^1] Which measure the distance of a distribution to uniformity; see e.g.
+We write $\Delta^{d}f$ for the map $(x,\mathbf{v})\mapsto\Delta^{d}_{\mathbf{v}}f(x)$.
+A function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ can be written [^1] Or: the unique reduced polynomial (variable degrees below $|\mathbb{F}|$) agreeing with $f$ has (total) degree below $d$.
+as a polynomial of degree below $d$ if and only if $\Delta^{d}f=0$ (i.e., $\Delta^{d}_{\mathbf{v}}f(x)=0$ for every $(x,\mathbf{v})\in(\mathbb{F}^{n})^{d+1}$).
+An important (qualitative) result about polynomials over $\mathbb{F}$, of characteristic $\operatorname{ch}(\mathbb{F})>d$, is that the derivative characterization is robust: $f$ is correlated with some polynomial of degree below $d$ if and only if $\Delta^{d}f$ is biased.
+This lies at the core of the theory of Gowers norms [Gow01, Tao12, HHL19], and is usually formulated using exponential sums: [^2] Which are closely related to the distance of a distribution from uniformity; see e.g.
 Claim 33 in [BV10].
@@
 where $e(y):=e^{2\pi iy/p}\colon\mathbb{F}\to\mathbb{C}$ and $\operatorname*{\mathbb{E}}_{x\in A}:=\frac{1}{|A|}\sum_{x\in A}$.
-We note that $\operatorname{bias}(f)\in[0,1]$, [^2] For example, if $\deg(f)<d$ then $\operatorname{bias}(\Delta^{d}f)=1$, while if $f$ is uniformly random then $\operatorname{bias}(\Delta^{d}f)=o(1)$.
-and that, explicitly, $\operatorname{bias}(\Delta^{d}f):=\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n},\mathbf{v}\in(\mathbb{F}^{n})^{d}}e(\Delta^{d}_{\mathbf{v}}f(x))$.
-The forward direction of this characterization—correlation implies bias—is not hard to prove, even quantitatively (e.g., using the Cauchy-Schwarz-Gowers inequality):
+We note that $\operatorname{bias}(f)\in[0,1]$, [^3] For example, if $\deg(f)<d$ then $\operatorname{bias}(\Delta^{d}f)=1$, while if $f$ is random then $\operatorname{bias}(\Delta^{d}f)=o(1)$ as $n\to\infty$.
+and that, explicitly, $\operatorname{bias}(\Delta^{d}f)=\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n},\mathbf{v}\in(\mathbb{F}^{n})^{d}}e(\Delta^{d}_{\mathbf{v}}f(x))$.
+[^4] The absolute value is not needed for $\Delta^{d}f$.
+The forward direction of this characterization—correlation implies bias—is not hard to prove, even quantitatively (e.g., using [^5] Indeed, if $\deg(P)<d$ then $\operatorname{bias}(\Delta^{d}f)=\operatorname{bias}(\Delta^{d}(f-P))\geq|\operatorname{bias}(f-P)|^{2^{d}}=\operatorname{cor}(f,P)^{2^{d}}$.
+the Cauchy-Schwarz-Gowers inequality or the monotonicity of Gowers norms [Tao12]):
 $\operatorname{bias}(\Delta^{d}f)\geq\operatorname{cor}_{<d}(f)^{2^{d}}$ (1)
@@
 The inverse direction is much harder.
-For $d=2$, the question falls in the domain of Fourier analysis, as $\operatorname{bias}(\Delta^{2}f)=\sum_{\zeta}|\hat{f}(\zeta)|^{4}$, and Parseval’s identity implies $\operatorname{cor}_{<2}(f)\geq\operatorname{bias}(\Delta^{2}f)^{1/2}$.
-However, standard Fourier analysis is insufficient for general $d$, and the inverse direction is only known to hold with an extremely weak bound [GM20].
-The inverse direction is often stated in terms of the Gowers (uniformity) norms of a complex-valued function.
+For $d=2$, the question falls in the domain of Fourier analysis, and it follows using Parseval’s identity that $\operatorname{cor}_{<2}(f)\geq\operatorname{bias}(\Delta^{2}f)^{1/2}$.
+[^6] Indeed, the magnitudes of the Fourier coefficients of $g=e(f)$ are $\operatorname{cor}(f,\ell)$, for any linear form $\ell$ (or character), and $\operatorname{bias}(\Delta^{2}f)=\sum_{\gamma}|\hat{g}(\gamma)|^{4}=\sum_{\ell}\operatorname{cor}(f,\ell)^{4}\leq\operatorname{cor}_{<2}(f)^{2}\sum_{\ell}\operatorname{cor}(f,\ell)^{2}=\operatorname{cor}_{<2}(f)^{2}$.
+For general $d$, however, the inverse question falls in the domain of higher-order Fourier analysis, and is only known to hold with a very weak bound [GM20] (stronger bounds are known in the integer setting [LSS24]).
+The inverse question is often stated in terms of the Gowers (uniformity) norms of a complex-valued function.
 For $g\colon\mathbb{F}^{n}\to\mathbb{C}$, the $U^{d}$-norm is $\lVert{g}\rVert_{U^{d}}=(\operatorname*{\mathbb{E}}_{(x,\mathbf{v})\in(\mathbb{F}^{n})^{d+1}}{\Delta^{*}}_{v^{(d)}}\cdots{\Delta^{*}}_{v^{(1)}}g(x))^{1/2^{d}}$, with $\Delta^{*}_{v}g(x)=g(x+v)\overline{g(x)}$, and the weak Gowers norm is $\lVert{g}\rVert_{u^{d}}=\max_{\deg(P)<d}\big|\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n}}g(x)e(-P(x))\big|$.
-Theorem 1.1 (Gowers inverse over finite fields [GT09, BTZ10, TZ10, TZ12]).
+Theorem 1.1 (Gowers inverse over finite fields [BTZ10, TZ10, TZ12]).
 For any $g\colon\mathbb{F}^{n}\to\mathbb{C}$ with $\|g\|_{\infty}\leq 1$,
@@
 $\exists\Phi>0\colon\operatorname{cor}_{<d}(f)\geq\Phi(\operatorname{bias}(\Delta^{d}f)).$ (2)
-We note that if the field has small characteristic $\operatorname{ch}(\mathbb{F})<d-1$, the maximum in Theorem 1.1, over $P$ with $\deg(P)<d$, must also range over so-called non-classical polynomials of degree below $d$ ([BS+22, GT09, LMS11]), which are functions $\mathbb{F}^{n}\to\mathbb{R}/\mathbb{Z}$ that involve division by powers of $\operatorname{ch}(\mathbb{F})$.
+We note that if the field has small characteristic $\operatorname{ch}(\mathbb{F})<d-1$, the maximum in Theorem 1.1, over $P$ with $\deg(P)<d$, must also range over so-called non-classical polynomials of degree below $d$ (see [TZ12] and [BS+22, GT09, LMS11]), which are functions $\mathbb{F}^{n}\to\mathbb{R}/\mathbb{Z}$ that involve division by powers of $\operatorname{ch}(\mathbb{F})$.
 In this paper we restrict ourselves to the high characteristic case (and so, to classical polynomials).
-The Polynomial Gowers Inverse conjecture posits that, in fact, the relationship between bias and correlation is about as tight as it could be; namely, that $\Phi(x)$ in Theorem 1.1 can be as large as a power of $x$ for any fixed $d$: $\Phi(x)\geq x^{O_{d}(1)}$ (this matches the easy direction $\Phi(x)\leq x^{2^{d}}$ in (1) up to the constant).
-It is perhaps the central conjecture of higher-order Fourier analysis, over finite fields and over the integers (e.g.
-[LSS24]), with applications in countless fields (see, e.g., [HHL19, Tao12, TZ10] and the references within).
+The Gowers inverse question goes back to foundational work of Gowers [Gow01], which led to the development of higher-order Fourier analysis, and seminal works of Green and Tao on additive patterns in the primes [GT10] and the $U^{3}$-norm [GT08], and of Samorodnitsky [Sam07] on polynomiality testing.
+It has found applications in a large number of areas (see, e.g., [BV10, HHL19, Pel24, Tao12, TZ10, Wol15] and the references within).
+However, the lack of strong bounds for Theorem 1.1 implies the same for the applications.
+The Polynomial Gowers Inverse conjecture [GT08] posits that, in fact, the relationship between bias and correlation is about as tight as it could be.
+Namely, the conjecture is that $\Phi(x)$ in Theorem 1.1 can be as large as a power of $x$ for any fixed $d$: $\Phi(x)\geq x^{O_{d}(1)}$ (this matches the easy direction $\Phi(x)\leq x^{1/2^{d}}$ in (1) up to the constant).
+It is perhaps the central conjecture of higher-order Fourier analysis.
 While a recent breakthrough in [GG+24, GG+25] was achieved in the case $d=3$, the Polynomial Gowers Inverse conjecture for any $d>3$ remains wide open.
@@
 In this paper we obtain an almost-polynomial bound for $\operatorname{GI}_{d}(d+1)$, the next open case.
-With a slight abuse of notation, we write $x^{\tilde{O}_{d}(1)}$ for $\exp(-O_{d}(\log(1/x)^{1+o(1)}))$.
+With a slight abuse of notation, given $x\in(0,1]$ we write $x^{\tilde{O}_{d}(1)}$ for $\exp(-O_{d}(\log(1/x)^{1+o(1)}))$.
+[^9] For completeness, set $0^{\tilde{O}_{d}(1)}=0$.
 (This is even stronger than a quasi-polynomial bound, which has a constant $O_{d}(1)$ at the top instead of 1+o(1)1+o(1).) [^10] More precisely, we get $\log(1/x)\cdot\log(\log(1/x)+1)$ in the exponent.
-(In fact, we even get a bound that improves as $|\mathbb{F}|$ increases, of the rough shape $x^{O_{d}(\log_{|\mathbb{F}|}(\log_{|\mathbb{F}|}(1/x)))}$, which for $|\mathbb{F}|\gg\log(1/x)$ is truly polynomial $x^{O_{d}(1)}$; see [MZ22].)
+(We even get a bound that improves as $|\mathbb{F}|$ increases, of the shape $x^{O_{d}(\log_{|\mathbb{F}|}(\log_{|\mathbb{F}|}(1/x)))}$, which for $|\mathbb{F}|\gg\log(1/x)$ is truly polynomial $x^{O_{d}(1)}$; see [MZ22].)
 Theorem 1.
@@
 Put differently, $\lVert{g}\rVert_{u^{d}}\geq\|g\|_{U^{d}}^{\tilde{O}_{d}(1)}$ for any polynomial phase $g\colon\mathbb{F}^{n}\to\mathbb{C}$ of degree at most $d+1$ (that is, $g=e(f)$ with $\deg(f)\leq d+1$).
-The main difficulty in the proof of Theorem 1 stems from the non-homogeneity of $f$.
-Standard proofs of the degree-$d$ case proceed by finding a decomposition of $f\in\operatorname{Poly}_{d}$ into lower-degree polynomials of length bounded in terms of $\operatorname{bias}(\Delta^{d}f)$, which is then used to obtain a significant correlation with a lower-degree polynomial.
-However, when $\deg(f)=d+1$, one obtains in this way only a decomposition of the top-degree part of $f$, and indeed, it is entirely possible for the degree-$d$ part $f_{d}$ of $f$ to not have any short decompositions despite having large $\operatorname{bias}(\Delta^{d}f)$ (e.g., $f=(x_{1}+1)h$ for a random $h$, see Remark 6.2).
-Furthermore, even if one were to successfully obtain a short decomposition of $f$ in terms of polynomials of degrees below $\deg(f)$, it is not clear how to use this to obtain a correlation with a polynomial of degree as small as $d-1=\deg(f)-2$; especially given that, unlike in the degree-$d$ case, one cannot guarantee a short decomposition with homogeneous factors.
-[^5] Concretely, the degree-$1$ factors in a decomposition of $f$ might not have a common zero (as they need not be homogeneous; e.g., $x_{1}$ and $x_{1}+1$), and so there might not even be a large subspace, or affine subspace, where the degree-$1$ factors can be discarded (which would have meant that all factors have degree at most $\deg(f)-2$).
-In the homogeneous setting, we can do more. Namely, we obtain almost-polynomial bounds for homogeneous polynomials of degree up to $2d$.
+A main difficulty in the proof of Theorem 1 stems from the non-homogeneity of $f$.
+Standard proofs of the degree-$d$ case utilize the fact that if $\deg(f)\leq d$, then the $d$-times derivative $\Delta^{d}f$ is a multilinear form, a class of polynomials for which we have good understanding of the interplay between structure and randomness.
+Concretely, we know how to convert high $\operatorname{bias}(\Delta^{d}f)$ into a decomposition of $f_{d}$—the top-degree homogeneous component of ff—in terms of few lower-degree polynomials.
+A standard pigeonhole argument is then used to obtain a significant correlation for $f_{d}$ with a polynomial $P$ of lower degree, $\deg(P)\leq d-1$, and thus for $f$ itself by absorbing its lower-degree components into $P$.
+However, when $\deg(f)=d+1$ and $f$ is not homogeneous, one obtains from the multilinear argument (via $\Delta^{d+1}f$, which is also biased) only a decomposition of the top-degree component $f_{d+1}$ of $f$, so absorbing the rest of ff—which is of degree dd—into a correlating polynomial $P$ with $\deg(P)\leq d-1$ is no longer an option.
+In fact, it is entirely possible for the degree-$d$ component $f_{d}$ of $f$ to not have any short decompositions, even if $\operatorname{bias}(\Delta^{d}f)$ is large (e.g., $f=(x_{1}+1)h$ for a random $h$, see Remark 6.2).
+Furthermore, even if one were to successfully obtain a short decomposition of $f$ itself in terms of polynomials of degrees below $\deg(f)$, it is not clear how to use this to obtain a correlation with a polynomial of degree as small as $d-1=\deg(f)-2$.
+[^11] Moreover, unlike in the degree-$d$ case, one cannot guarantee a short decomposition with homogeneous factors.
+In particular, the degree-$1$ factors in a decomposition of $f$ might not have a common zero (as they need not be homogeneous; e.g., $x_{1}$ and $x_{1}+1$), and so there might not even be a large (affine) subspace where the degree-$1$ factors can be discarded (which would have at least meant that all factors have degree at most $\deg(f)-2$ there).
+In the homogeneous setting, we can do more. Namely, we obtain almost-polynomial bounds for homogeneous polynomials of degree smaller than $2d$.
 Theorem 2.
 Let $d\geq 1$. For homogeneous $f\colon\mathbb{F}^{n}\to\mathbb{F}$ with $\deg(f)<2d$, where $\deg(f)<\operatorname{ch}(\mathbb{F})$,
-$\operatorname{cor}_{<d}(f)\geq\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}$
+$\operatorname{cor}_{<d}(f)\geq\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}.$
 The proof of Theorem 2 relies on a new, general correlation lemma.
-Unlike standard correlation lemmas in the literature (e.g.
-[Jan19], proof of Theorem 1.6, or [GT09, CM23] and many more) which require a full decomposition of the given polynomial, ours requires a more relaxed condition, closely related to membership in an appropriate ideal.
+Standard correlation lemmas/arguments in the literature (e.g.
+[Jan20], proof of Theorem 1.6; or [GT09], beginning of the proof of Theorem 1.4; or [KL08], proof of Theorem 3; or [TZ12], beginning of the proof of Theorem 1.10 from Theorem 1.18; and many more) use a simple pigeonhole-type argument on Fourier coefficients to obtain significant correlation with a lower degree polynomial, provided we have a full decomposition of the given polynomial.
+Our new correlation lemma, on the other hand, requires a more relaxed condition, closely related to membership in an appropriate ideal.
 Furthermore, as another application of our new correlation lemma, we show in Section 3.1 how it can be used to provide an alternative proof of a polynomial bound for $\operatorname{GI}_{d}(d)$ from a weak version of the partition-versus-analytic rank dichotomy [Mil25].
@@
 2.1 Preliminaries
-Henceforth, $\mathbb{F}$ is any finite field.
+From now on $\mathbb{F}$ is an arbitrary finite field.
 We write $\operatorname{\mathfrak{bias}}$ (in Fraktur letters) for the bias as a complex number: for any function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ over a finite field $\mathbb{F}$, denote
@@
 We denote the indicator function for a set $S$ by $\mathbf{1}_{S}(x)=\begin{cases}1&x\in S\\ 0&x\notin S\end{cases}$.
+We denote by $\operatorname{Poly}_{k}(\mathbb{F})$ the set of (formal) polynomials of degree at most $k$, over the field $\mathbb{F}$, in any (finite) number of variables $x_{1},\ldots,x_{n}$.
+[^12] When combining polynomials, they are understood to be defined on the same variable set (adding dummy variables if needed).
 A form is a homogeneous polynomial.
-We denote by $\operatorname{Form}_{k}(\mathbb{F})$ the set of forms of degree $k$, over the field $\mathbb{F}$, in any (finite) number of variables, or simply $\operatorname{Form}_{k}$ when $\mathbb{F}$ is immaterial.
-Similarly, we denote by $\operatorname{Poly}_{k}(\mathbb{F})$ the set of polynomials of degree at most $k$ over $\mathbb{F}$ is any number of variables.
+We denote by $\operatorname{Form}_{k}(\mathbb{F})$ the set of forms of degree $k$ ($\geq 0$) over $\mathbb{F}$ in any number of variables.
+We simply write $\operatorname{Form}_{k}$ when $\mathbb{F}$ is immaterial, or $\operatorname{Form}$ when $k$ is immaterial as well.
 Note that any $f\in\operatorname{Poly}_{k}$ is the sum of forms, namely, the homogeneous components of $f$ of each degree: $f=\sum_{i=0}^{k}f_{i}$ with $f_{i}\in\operatorname{Form}_{i}$.
@@
 $f=\sum_{i=1}^{r}g_{i}h_{i}$
-for some forms $g_{i}$ and $h_{i}$ of degree below $d$.
-A form is multilinear if it is linear in each of its variable vectors; more concretely, a multilinear form of degree $d$, or a $d$-linear form for short, is of the form $T=T(x^{(1)},\ldots,x^{(d)}):(\mathbb{F}^{n})^{d}\to\mathbb{F}$ and is linear in each of $x^{(1)},\ldots,x^{(d)}$.
-Equivalently, any monomial in the support of $T$ has one variable from each variable vector.
+for some forms $g_{i}$ and $h_{i}$ of (positive) degree smaller than $\deg(f)$.
+A form is multilinear if it is linear in each of its variable vectors; more concretely, a multilinear form of degree $d$ ($\geq 1$), or a $d$-linear form for short, is of the form $T=T(x^{(1)},\ldots,x^{(d)}):(\mathbb{F}^{n})^{d}\to\mathbb{F}$ and is linear in each of $x^{(1)},\ldots,x^{(d)}$.
+Equivalently, any monomial in the support of $T$ has one variable from each of $x^{(1)},\ldots,x^{(d)}$.
 The partition rank [Nas20] of a $d$-linear form $T$, denoted by $\operatorname{PR}(T)$, is the minimal $r$ for which $T$ is the sum of reducible $d$-linear forms:
 $T=\sum_{i=1}^{r}R_{i}Q_{i}$
-for some $R_{i}=R_{i}(x_{I_{i}})$ and $Q_{i}=Q_{i}(x_{I_{i}^{c}})$ with $\emptyset\subsetneq I_{i}\subsetneq[d]$, where for $x\in(\mathbb{F}^{n})^{d}$ we write $x_{I}$ for the restriction of $x$ to the coordinates corresponding to $I\subseteq[d]$.
+for some multilinear forms $R_{i}=R_{i}(x_{I_{i}})$ and $Q_{i}=Q_{i}(x_{I_{i}^{c}})$ with $\emptyset\subsetneq I_{i}\subsetneq[d]$, where for $x\in(\mathbb{F}^{n})^{d}$ we write $x_{I}$ for the restriction of $x$ to the coordinates corresponding to $I\subseteq[d]$.
 A related notion to the partition rank is that of the analytic rank [GW11] over a finite field $\mathbb{F}$, denoted by $\operatorname{AR}(T)$:
 $\operatorname{AR}(T):=-\log_{|\mathbb{F}|}(\operatorname{bias}(T)).$
-We note that for a multilinear $T$, we have $\operatorname{bias}(T)=\mathbb{P}(T=0)-\mathbb{P}(T=c)$ for any nonzero $c\in\mathbb{F}$.
-The polarization of a form $g\in\operatorname{Form}_{k}$, which we denote by $\overline{g}$, is the unique (degree-$k$) multilinear form satisfying $\overline{g}(x,\ldots,x)=k!g(x)$.
-(Note that if $g\colon\mathbb{F}^{n}\to\mathbb{F}$ then $\overline{g}\colon(\mathbb{F}^{n})^{k}\to\mathbb{F}$.) Equivalently, $\overline{g}$ can be obtained as $\overline{g}=\Delta^{k}g$, the discrete derivative of order $k$; that is, $\overline{g}(\mathbf{x})=\Delta^{k}_{\mathbf{x}}g$ where $\mathbf{x}=(x^{(1)},\ldots,x^{(k)})$.
-(See Fact 5.7 below.) Note that $\Delta^{k}_{\mathbf{x}}g$ depends only on $\mathbf{x}$, being independent of the point at which we take the discrete derivative (as $k=\deg(g)$).
-For example, for $g(x)=x_{1}x_{2}$, we have $\overline{g}(x,y)=x_{1}y_{2}+y_{1}x_{2}$, and for $g(x)=x_{1}^{2}$ we have $\overline{g}(x,y)=2x_{1}y_{1}$.
-We denote by $\partial_{c}f$ the formal derivative of a polynomial $f$ in direction $c$, and denote by $\boldsymbol{\nabla}f$ the formal gradient of $f$, so that $\partial_{c}f(x)=c\cdot\boldsymbol{\nabla}f$.
-Note that the formal derivative is well defined for polynomials over any field; by linearity, it suffices to define it for monomials: for $f=\prod_{i}x_{i}^{d_{i}}$, $\frac{\partial f}{\partial x_{j}}:=d_{j}x_{j}^{d_{j}-1}\prod_{i\neq j}x^{d_{i}}$, $\boldsymbol{\nabla}f:=(\frac{\partial f}{\partial x_{i}})_{i=1}^{n}$ and $\partial_{c}f=c\cdot\boldsymbol{\nabla}f=\sum_{i=1}^{n}c_{i}\frac{\partial f}{\partial x_{i}}$.
-A fundamental part of our argument relies on the best bounds known for the rank of biased homogeneous polynomials (see, e.g., Corollary 8.5 in [MZ22] [^6] One can naively extend the definition of $\operatorname{rk}$ to non-homogeneous polynomials $f\in\operatorname{Poly}_{k}$ by defining $\operatorname{rk}(f):=\operatorname{rk}(f_{k})$, but this is not useful in this paper as this does not give a decomposition of $f$, only of $f_{k}$.).
+We note that for $T$ multilinear, $\operatorname{bias}(T)$ also has an elementary definition: $\operatorname{bias}(T)=\mathbb{P}(T=0)-\mathbb{P}(T=c)$ ($\geq 0$), for any choice of nonzero $c\in\mathbb{F}$.
+[^13] Indeed, multilinearity implies that $\mathbb{P}(T=t)$ is the same for every $t\in\mathbb{F}^{\times}$, so $\operatorname{\mathfrak{bias}}(T)=\sum_{t\in\mathbb{F}}\mathbb{P}(T=t)\chi(t)=\mathbb{P}(T=0)+\mathbb{P}(T=c)(\sum_{t\in\mathbb{F}}\chi(t)-\chi(0))=\mathbb{P}(T=0)-\mathbb{P}(T=c)$.
+Moreover, $\operatorname{AR}(T)\leq\operatorname{PR}(T)$ (e.g., [Lov19]).
+A fundamental part of our argument relies on the best bound known for the partition rank of biased multilinear forms.
+[^14] For completeness, note that if $d=1$ then $\operatorname{AR}(T)=\operatorname{PR}(T)=\infty$.
 We write $\tilde{O}_{d}(x):=O_{d}(x^{1+o(1)})$.
@@
 Theorem 2.1 ([MZ22]).
-For every form $g\in\operatorname{Form}_{d}$, over any finite field $\mathbb{F}$ with $\operatorname{ch}(\mathbb{F})>d$,
-$\operatorname{rk}(g)\leq\tilde{O}_{d}(\operatorname{AR}(g)).$
-Moreover, for every $d$-linear form $T$,
-$\operatorname{PR}(T)=\tilde{O}_{d}(\operatorname{AR}(T))$
+For every $d$-linear form $T$ over any finite field,
+$\operatorname{PR}(T)=\tilde{O}_{d}(\operatorname{AR}(T)).$
+The polarization of a form $g\in\operatorname{Form}_{k}$, which we denote by $\overline{g}$, is the $k$-linear form $\overline{g}=\Delta^{k}g$, given by the discrete derivative of order $k$; that is, $\overline{g}(\mathbf{x})=\Delta^{k}_{\mathbf{x}}g$ where $\mathbf{x}=(x^{(1)},\ldots,x^{(k)})$.
+(Note that if $g\colon\mathbb{F}^{n}\to\mathbb{F}$ then $\overline{g}\colon(\mathbb{F}^{n})^{k}\to\mathbb{F}$; indeed, $\Delta^{k}_{\mathbf{x}}g$ depends only on $\mathbf{x}$, being independent of the point at which we take the discrete derivative, as $k=\deg(g)$.) If $k<\operatorname{ch}(\mathbb{F})$, $g$ can be recovered from its polarization: $g(x)=\overline{g}(x,\ldots,x)/k!$ (and $\overline{g}$ is the unique symmetric $k$-linear form satisfying this); see Fact 5.7 below.
+For example, for $g(x)=x_{1}x_{2}$, we have $\overline{g}(x,y)=x_{1}y_{2}+y_{1}x_{2}$, and for $g(x)=x_{1}^{2}$ we have $\overline{g}(x,y)=2x_{1}y_{1}$ (so $\overline{g}(x,x)$ fails to recover $g$ only when $\operatorname{ch}(\mathbb{F})=2$).
+We denote by $\partial_{c}f$ the formal derivative of a polynomial $f$ in direction $c$, and denote by $\boldsymbol{\nabla}f$ the formal gradient of $f$, so that $\partial_{c}f(x)=c\cdot\boldsymbol{\nabla}f$.
+Note that the formal derivative is well defined for polynomials over any field; by linearity, it suffices to define it for monomials: for $f=\prod_{i}x_{i}^{d_{i}}$, $\frac{\partial f}{\partial x_{j}}:=d_{j}x_{j}^{d_{j}-1}\prod_{i\neq j}x^{d_{i}}$, $\boldsymbol{\nabla}f:=(\frac{\partial f}{\partial x_{i}})_{i=1}^{n}$ and $\partial_{c}f=c\cdot\boldsymbol{\nabla}f=\sum_{i=1}^{n}c_{i}\frac{\partial f}{\partial x_{i}}$.
 2.2 Proof outline
@@
 Call that bias $\delta$.
-Using a partition-versus-analytic rank bound, such as Theorem 2.1, one obtains a partition-rank decomposition for $\Delta^{d}f$ of length $\log(\frac{1}{\delta})^{\tilde{O}_{d}(1)}$.
-By plugging in the diagonal $(x,\ldots,x)$ and dividing by $d!$, we obtain a decomposition of $f_{d}$ (the degree-$d$ part of $f$), meaning $\operatorname{rk}(f_{d})\leq\log(\frac{1}{\delta})^{\tilde{O}_{d}(1)}$.
-Writing $f=\sum_{i}\alpha_{i}\beta_{i}$, a standard Fourier/pigeonhole-type argument gives a correlation for $f_{d}$ of at least $\delta^{\tilde{O}_{d}(1)}$ with some member of the span of the $\alpha_{i}$ and $\beta_{i}$—which is of degree at most d1d-1—and thus a correlation for $f$ itself with such a lower-degree polynomial.
-The proof of Theorem 2, for homogeneous polynomials, follows this classical recipe, except it utilizes a new correlation lemma (Lemma 3.1) to obtain a correlation with a linear combination of only the lower-degree factors from each product $\alpha_{i}\beta_{i}$ in the decomposition, which is a polynomial of degree as low as $\deg(f)/2$.
+Using a partition-versus-analytic rank bound, such as Theorem 2.1, one obtains a partition-rank decomposition for $\Delta^{d}f$ of length $\tilde{O}_{d}(\log_{|\mathbb{F}|}(1/\delta))$.
+By plugging in the diagonal $(x,\ldots,x)$ and dividing by $d!$, we obtain a decomposition of $f_{d}$ (the degree-$d$ component of $f$), meaning $\operatorname{rk}(f_{d})\leq\tilde{O}_{d}(\log_{|\mathbb{F}|}(1/\delta))$.
+Writing $f_{d}=\sum_{i}\alpha_{i}\beta_{i}$, a standard Fourier/pigeonhole-type argument gives a correlation for $f_{d}$ of at least $\delta^{\tilde{O}_{d}(1)}$ with some linear combination of $\alpha_{i}$ and $\beta_{i}$—which is of degree at most d1d-1—and thus a correlation for $f$ itself with such a lower-degree polynomial.
+The proof of Theorem 2, for homogeneous polynomials, follows this classical recipe, except it utilizes a new correlation lemma (Lemma 3.1) to obtain a correlation with a linear combination of only the lower-degree factors from each product $\alpha_{i}\beta_{i}$ in the decomposition, which is a polynomial of degree at most $\deg(f)/2$ (<d<d).
 The difficulty in the non-homogeneous case of Theorem 1 is twofold.
@@
 The objective here was to find a notion of rank for which the following holds: whenever $\operatorname{rk}^{*}(f)$ is small, one can find a large correlation of $f$ with a polynomial of degree at most $\deg(f)-2$ (whereas for the usual notion of rank, one can only generally get $\deg(f)-1$).
-It turns out that the requirement needed is quite weak: if $f=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ is a $\operatorname{rk}$-decomposition, for it to be a $\operatorname{rk}^{*}$ decomposition, we require the degree-$1$ polynomials in the decomposition to have linearly independent linear components.
-This requirement guarantees that the zero-locus of these polynomials is fairly large, which is essential for us to obtain significant correlation.
+It turns out that the requirement needed is quite weak: for a decomposition $f=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ ($1\leq\deg(\alpha_{i})\leq\deg(\beta_{i})<\deg(f)$) to be a $\operatorname{rk}^{*}$-decomposition, we require the degree-$1$ polynomials $\alpha_{i}$ (in fact, only those with $\deg(\alpha_{i}\beta_{i})=\deg(f)$) to have linearly independent linear components.
+This requirement guarantees that the zero-locus of these polynomials is fairly large after an appropriate perturbation (Lemma 4.8), which is essential for us to obtain significant correlation.
 We now explain how the $\operatorname{rk}^{*}$-decomposition of $f$ is obtained.
-Write $g=f_{d+1}$ and $h=f_{d}$ for the homogeneous parts of degree $d+1$ and $d$ respectively.
-Proposition 5.1 provides a simple formula for $\Delta^{d}f$ for any polynomial $f$ of degree $d+1$.
-Proposition 6.4 shows that decomposing $g$ by the standard polarization process goes smoothly, thanks to the chain rule for bias (Lemma 6.3), which tells us that $\Delta^{d}g$} is biased.
+Write $g=f_{d+1}$ and $h=f_{d}$ for the homogeneous components of degree $d+1$ and $d$ respectively.
+Proposition 5.1 provides a simple formula for $\Delta^{d}g$.
+Proposition 6.4 in particular shows that decomposing $g$ by the standard polarization process goes smoothly, thanks to a chain rule for bias (Lemma 6.3).
 The decomposition of $h$ is a more delicate issue, since the polarization $\overline{h}$ need not be biased.
-However, we prove that $\overline{h}$ is biased when adding an appropriate partial derivative $\partial_{c}\overline{g}$; moreover, the sum $\overline{h}+\partial_{c}\overline{g}$ is a $d$-linear form, hence another application of Theorem 2.1 gives a decomposition to this sum.
-To deal with the summand $\partial_{c}\overline{g}$ we added, we use a chain rule for $\operatorname{rk}^{*}$ (Proposition 4.5).
+However, we prove that $\overline{h}$ is biased when adding an appropriate partial derivative $\partial_{c}\overline{g}$ (with respect to one of the $d$ vector variables); moreover, the sum $\overline{h}+\partial_{c}\overline{g}$ is a $d$-linear form, hence an application of Theorem 2.1 gives a decomposition to this sum.
+To deal with the summand $\partial_{c}\overline{g}$ we added, we use a chain rule for $\operatorname{rk}^{*}$ (Proposition 4.5) and the fact that polarization and formal differentiation commute (Lemma 5.2).
 All of the above allows us to obtain a non-homogeneous decomposition for $f$, which is nevertheless useful for us since it is a short $\operatorname{rk}^{*}$-decomposition.
 Using the control we have over its linear factors, Proposition 4.7 allows us then to use this special decomposition to obtain significant correlation with a polynomial of degree $\deg(f)-2\leq d-1$, as needed.
-We finally remark that we cannot utilize the arguments leading to the polynomial bound for $\operatorname{GI}_{d}(d)$, in Theorem 3.4, in order to obtain a similar bound for degree $d+1$, since our argument requires more control over the factors than is afforded by that Theorem.
+We finally remark that we cannot utilize the arguments leading to the polynomial bound for $\operatorname{GI}_{d}(d)$, in Theorem 3.4, in order to obtain a similar bound for degree $d+1$, since the latter does not help in providing a decomposition, which our arguments require.
 3 Correlation lemma, the homogeneous and the degree-$d$ cases
@@
 For a tuple of functions $\mathbf{A}=(A_{1},\ldots,A_{m})$, their joint zero set is $\operatorname{Z}(\mathbf{A})=\{x\,|\,\forall i\colon A_{i}(x)=0\}$.
-For an event $E$, we abbreviate $\operatorname{\mathfrak{bias}}(g\,|\,E)=\operatorname*{\mathbb{E}}[\chi(g)\,|\,E]$.
+For an event $E$, we abbreviate $\operatorname{\mathfrak{bias}}(f\,|\,E)=\operatorname*{\mathbb{E}}[\chi(f)\,|\,E]$.
 Lemma 3.1 (average correlation lemma).
-Let $g,A_{1},\ldots,A_{m}\colon\mathbb{F}^{n}\to\mathbb{F}$ be any functions, and put $\mathbf{A}=(A_{1},\ldots,A_{m})$. Then
-$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\cdot\operatorname{\mathfrak{bias}}(g\,|\,\operatorname{Z}(\mathbf{A})).$
+Let $f,A_{1},\ldots,A_{m}\colon\mathbb{F}^{n}\to\mathbb{F}$ be any functions, and put $\mathbf{A}=(A_{1},\ldots,A_{m})$. Then
+$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\cdot\operatorname{\mathfrak{bias}}(f\,|\,\operatorname{Z}(\mathbf{A})).$
+For completeness, if $\operatorname{Z}(\mathbf{A})=\emptyset$ then the right hand side should be interpreted as 0.
 Proof.
 By direct computation:
-$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)$ $=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\chi\Big(g(x)-\sum_{i=1}^{m}c_{i}A_{i}(x)\Big)$
-$=\operatorname*{\mathbb{E}}_{x}\chi(g(x))\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\chi\Big(\sum_{i=1}^{m}c_{i}A_{i}(x)\Big)$
-$=\operatorname*{\mathbb{E}}_{x}\chi(g(x))\mathbf{1}_{\operatorname{Z}(\mathbf{A})}(x)$
-$=|\mathbb{F}|^{-n}\sum_{x\in\operatorname{Z}(\mathbf{A})}\chi(g(x))$
-$=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\frac{1}{|\operatorname{Z}(\mathbf{A})|}\sum_{x\in\operatorname{Z}(\mathbf{A})}\chi(g(x))$
-$=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\operatorname{\mathfrak{bias}}(g\,|\,\operatorname{Z}(\mathbf{A})).\qed$
-We will also need the following Theorem due to Warning (for a short proof see [Asg23]).
+$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)$ $=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\chi\Big(f(x)-\sum_{i=1}^{m}c_{i}A_{i}(x)\Big)$
+$=\operatorname*{\mathbb{E}}_{x}\chi(f(x))\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\chi\Big(\sum_{i=1}^{m}c_{i}A_{i}(x)\Big)$
+$=\operatorname*{\mathbb{E}}_{x}\chi(f(x))\mathbf{1}_{\operatorname{Z}(\mathbf{A})}(x)$
+$=|\mathbb{F}|^{-n}\sum_{x\in\operatorname{Z}(\mathbf{A})}\chi(f(x)),$
+where the third equality is the fundamental fact that $\operatorname*{\mathbb{E}}_{c}\chi(c\cdot a)$ is $1$ if $a=0$, and otherwise 0 (equivalently, the orthogonality of additive characters).
+If $\operatorname{Z}(\mathbf{A})\neq\emptyset$, we therefore get
+$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\frac{1}{|\operatorname{Z}(\mathbf{A})|}\sum_{x\in\operatorname{Z}(\mathbf{A})}\chi(f(x))\\ =\mathbb{P}(\operatorname{Z}(\mathbf{A}))\operatorname{\mathfrak{bias}}(f\,|\,\operatorname{Z}(\mathbf{A})).\qed$
+We will also need the following Theorem due to Warning (for a short proof see [Asg18]).
 Theorem 3.2 (Warning’s second theorem [War35]).
 For all polynomials $f_{1},\ldots,f_{m}\colon\mathbb{F}^{n}\to\mathbb{F}$ over a finite field $\mathbb{F}$, if $\operatorname{Z}(f_{1},\ldots,f_{m})\neq\emptyset$ then $|\operatorname{Z}(f_{1},\ldots,f_{m})|\geq|\mathbb{F}|^{n-\sum_{i=1}^{m}\deg(f_{i})}$.
-We obtain the following correlation bound, where $\operatorname{cor}_{d}(g)=\max_{\deg(P)\leq d}\operatorname{cor}(g,P)$.
+We obtain the following correlation bound, which only makes an assumption on the zero set of $f$ (or on an ideal containing $f$).
+Recall the notation $\operatorname{cor}_{<d}(f)=\max_{\deg(P)<d}\operatorname{cor}(f,P)$.
 Corollary 3.3.
-If $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(g)$ for some $\mathbf{A}=(A_{1},\ldots,A_{m})\in\operatorname{Form}_{d}(\mathbb{F})^{m}$, then $\operatorname{cor}_{d}(g)\geq|\mathbb{F}|^{-dm}$.
+If $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(f)$ for some $\mathbf{A}=(A_{1},\ldots,A_{m})\in\operatorname{Form}(\mathbb{F})^{m}$ with $1\leq\deg(A_{i})<d$, then $\operatorname{cor}_{<d}(f)\geq|\mathbb{F}|^{-(d-1)m}$.
 Proof.
-Apply Lemma 3.1 to obtain
-$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{bias}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)$ $\geq\Big|\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)\Big|$
-$=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\cdot\operatorname{bias}(g\,|\,\operatorname{Z}(\mathbf{A}))=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\geq|\mathbb{F}|^{-dm},$
-where the last inequality applies Theorem 3.2, using that $0\in\operatorname{Z}(A_{1},\ldots,A_{m})$ by homogeneity.
-Thus, there exists a form $P=\sum_{i=1}^{m}c_{i}A_{i}$ of degree at most $d$ such that $\operatorname{bias}(g-P)\geq|\mathbb{F}|^{-dm}$.
+Put $D=\sum_{i=1}^{m}\deg(A_{i})$. Apply Lemma 3.1 to obtain
+$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{bias}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)$ $\geq\Big|\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)\Big|$
+[…]
+where the last inequality applies Warning’s second theorem (Theorem 3.2), using the fact that $0\in\operatorname{Z}(A_{1},\ldots,A_{m})$ by homogeneity.
+Thus, some $P=\sum_{i=1}^{m}c_{i}A_{i}\in\operatorname{Poly}_{d-1}$ satisfies $\operatorname{bias}(f-P)\geq|\mathbb{F}|^{-D}\geq|\mathbb{F}|^{-(d-1)m}$.
 ∎
@@
 Let $f:\mathbb{F}^{n}\to\mathbb{F}$ be a homogeneous polynomial with $k:=\deg(f)<2d$, where $\mathbb{F}=\mathbb{F}_{q}$ with $\operatorname{ch}(\mathbb{F}_{q})>k$, and put $\operatorname{bias}(\Delta^{d}f)=\delta$.
-First, we observe that $\operatorname{bias}(\overline{f})\geq\delta^{2^{d}}$.
-Indeed, using the monotonicity of Gowers norms, $\lVert{g}\rVert_{U^{k}}\geq\lVert{g}\rVert_{U^{d}}$ for $k\geq d$, and using the fact that, by definition, $\operatorname{bias}(\Delta^{t}f)=\lVert{e(f)}\rVert_{U^{t}}^{2^{t}}$ for any $t$, we have that
+We assume $k\geq d\geq 2$ and $\delta>0$, as otherwise there is nothing to prove.
+First, we observe that $\operatorname{bias}(\overline{f})\geq\delta^{2^{d}}$, recalling that $\overline{f}=\Delta^{k}f$ denotes polarization.
+Indeed, using the monotonicity of Gowers norms, $\lVert{g}\rVert_{U^{k}}\geq\lVert{g}\rVert_{U^{d}}$ for $k\geq d$ and any $g\colon\mathbb{F}^{n}\to\mathbb{C}$, and using the fact that, by definition, $\operatorname{bias}(\Delta^{t}f)=\lVert{e(f)}\rVert_{U^{t}}^{2^{t}}$ for any $t$, we have that
 $\operatorname{bias}(\overline{f})=\operatorname{bias}(\Delta^{k}f)=\lVert{e(f)}\rVert^{2^{k}}_{U^{k}}\geq\lVert{e(f)}\rVert^{2^{k}}_{U^{d}}=\operatorname{bias}(\Delta^{d}f)^{2^{k-d}}\geq\delta^{2^{d}}.$
@@
 $\overline{f}=\sum_{i=1}^{r}Q_{i}R_{i}$
-with $Q_{i},R_{i}$ multilinear forms of degrees at most k1k-1, and
-$r\leq\tilde{O}_{k}(\log_{p}(1/\operatorname{bias}(\overline{f})))=\tilde{O}_{d}(\log_{q}(1/\delta)).$
+with $Q_{i},R_{i}$ multilinear forms of degrees at most k1k-1, on disjoint sets of vector variables, and
+$r\leq\tilde{O}_{k}(\log_{q}(1/\operatorname{bias}(\overline{f})))=\tilde{O}_{d}(\log_{q}(1/\delta)).$
 Assume without loss of generality that $\deg(Q_{i})\leq\deg(R_{i})$ for every $i$.
 Since $\deg(Q_{i})+\deg(R_{i})=k$, we have $\deg(Q_{i})\leq k/2$.
-Thus, $\overline{f}\in\langle Q_{1},\ldots,Q_{r}\rangle$.
+Thus, we can put $\overline{f}$ in the ideal $\overline{f}\in\langle Q_{1},\ldots,Q_{r}\rangle$.
 Plug in the diagonal: let $A_{i}(x)=Q_{i}(x,\ldots,x)$, and observe that $f(x)=\frac{1}{k!}\overline{f}(x,\ldots,x)$, using the homogeneity of $f$ and $\operatorname{ch}(\mathbb{F})>k$.
 It follows that $f\in\langle A_{1},\ldots,A_{r}\rangle$ with each $A_{i}$ homogeneous of degree at most $k/2<d$, by assumption.
-We are now done by Corollary 3.3:
-$\operatorname{cor}_{<d}(f)\geq q^{-dr}\geq p^{-\tilde{O}_{d}(\log_{q}(1/\delta))}=\delta^{\tilde{O}_{d}(1)}.\qed$
+We may assume $\deg(A_{i})\geq 1$, as otherwise $A_{i}=0$ so we may discard it.
+Since $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(f)$, we are now done by Corollary 3.3:
+$\operatorname{cor}_{<d}(f)\geq q^{-dr}\geq q^{-\tilde{O}_{d}(\log_{q}(1/\delta))}=\delta^{\tilde{O}_{d}(1)}.\qed$
 3.1 A remark on the case of degree-$d$ polynomials
@@
 Let $f\colon\mathbb{F}^{n}\to\mathbb{F}$ with $\deg(f)\leq d$ and $\operatorname{bias}(\Delta^{d}f)=\delta$, where $\mathbb{F}=\mathbb{F}_{p}$ and $d<p$.
-Recall that $\Delta^{d}f\colon(\mathbb{F}^{n})^{d}\to\mathbb{F}$ is a multilinear form of degree $d$, and let us write $\Delta^{d}f(x^{(1)},\ldots,x^{(d)})=\sum_{i=1}^{n}x^{(d)}_{i}T_{i}(x^{(1)},\ldots,x^{(d-1)})$.
+We assume $\deg(f)=d\geq 2$ and $\delta>0$, as otherwise there is nothing to prove.
+Recall that $\Delta^{d}f\colon(\mathbb{F}^{n})^{d}\to\mathbb{F}$ is a $d$-linear form, and let us write $\Delta^{d}f(x^{(1)},\ldots,x^{(d)})=\sum_{i=1}^{n}x^{(d)}_{i}T_{i}(x^{(1)},\ldots,x^{(d-1)})$, so that each $T_{i}\colon(\mathbb{F}^{n})^{d-1}\to\mathbb{F}$ is a (d1)(d-1)-linear form.
 Put $\mathbf{T}=(T_{1},\ldots,T_{n})$, and consider the multilinear variety $V=\operatorname{Z}(\mathbf{T})\subseteq(\mathbb{F}^{n})^{d-1}$.
-Note that the density of $V$ is $\mathbb{P}(\mathbf{x}\in\operatorname{Z}(\mathbf{T}))=\operatorname{bias}(\Delta^{d}f)=\delta$.
-By Theorem 3.4, there are multilinear forms $\mathbf{S}=(S_{1},\ldots,S_{m})$, with each $S_{i}\colon(\mathbb{F}^{n})^{d-1}\to\mathbb{F}$ of degrees at most d1d-1, such that $\operatorname{Z}(\mathbf{S})\subseteq\operatorname{Z}(\mathbf{T})$ ($\subseteq\operatorname{Z}(\Delta^{d}f)$) and $m\leq O_{d}(\log_{p}(1/\delta)+1)$.
+Note that the density of $V$ is $\mathbb{P}_{\mathbf{x}\in(\mathbb{F}^{n})^{d-1}}(\mathbf{x}\in V)=\operatorname{bias}(\Delta^{d}f)=\delta$; indeed, this is the standard observation $\operatorname{bias}(\Delta^{d}f)=\operatorname*{\mathbb{E}}_{\mathbf{x}=(x^{(1)},\ldots,x^{(d-1)})}\operatorname*{\mathbb{E}}_{x^{(d)}}\chi(x^{(d)}\cdot\mathbf{T}(\mathbf{x}))=\operatorname*{\mathbb{E}}_{\mathbf{x}}\mathbf{1}_{V}(\mathbf{x})=\mathbb{P}_{\mathbf{x}}(\mathbf{x}\in V)$.
+By Theorem 3.4 with d1d-1, there are multilinear forms $\mathbf{S}=(S_{1},\ldots,S_{m})$, with each $S_{i}\colon(\mathbb{F}^{n})^{d-1}\to\mathbb{F}$ of degrees at most d1d-1, such that $\operatorname{Z}(\mathbf{S})\subseteq\operatorname{Z}(\mathbf{T})$ and $m\leq O_{d}(\log_{p}(1/\delta)+1)$.
 We observe that for any nonzero $d$-linear form $T$, $\operatorname{AR}(T)\geq 2^{-d}$; this follows for example from Claim 3.3 in [Lov19].
-Since we may assume $\Delta^{d}f\neq 0$, as otherwise $\delta=\operatorname{bias}(\Delta^{d}f)=1$ and there is nothing to prove, we have that $\log_{p}(1/\delta)=\operatorname{AR}(\Delta^{d}f)\geq 2^{-d}$.
+Since $\Delta^{d}f\neq 0$, as otherwise $\deg(f)<d$, we have that $\log_{p}(1/\delta)=\operatorname{AR}(\Delta^{d}f)\geq 2^{-d}$.
 Thus, $\log_{p}(1/\delta)+1\leq(2^{d}+1)\log_{p}(1/\delta)$, and therefore
 $m\leq O_{d}(\log_{p}(1/\delta)).$ (3)
-Plugging in the diagonal, let $A_{i}(x)=S_{i}(x,\ldots,x)$, homogeneous with $\deg(A_{i})\leq\det(S_{i})\leq d-1$.
+Plugging in the diagonal, let $A_{i}(x)=S_{i}(x,\ldots,x)$, homogeneous with $\deg(A_{i})\leq\deg(S_{i})\leq d-1$.
 Put $\mathbf{A}=(A_{1},\ldots,A_{m})$.
 We have $f_{d}(x)=(1/d!)\Delta^{d}f(x,\ldots,x)$ using $d<p$.
-Since $\operatorname{Z}(\mathbf{S})\subseteq\operatorname{Z}(\Delta^{d}f)$, we have that $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(f_{d})$.
-[^8] Indeed, $\operatorname{Z}(\mathbf{A})=\{x\,|\,\mathbf{A}(x)=0\}=\{x\,|\,\mathbf{S}(x,\ldots,x)=0\}\subseteq\{x\,|\,\Delta^{d}f(x,\ldots,x)=0\}=\operatorname{Z}(f_{d})$.
+Since $\operatorname{Z}(\mathbf{S})\times\mathbb{F}^{n}\subseteq\operatorname{Z}(\mathbf{T})\times\mathbb{F}^{n}\subseteq\operatorname{Z}(\Delta^{d}f)$, we have that $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(f_{d})$.
+[^16] Indeed, $\operatorname{Z}(\mathbf{A})=\{x\,|\,\mathbf{A}(x)=0\}=\{x\,|\,\mathbf{S}(x,\ldots,x)=0\}\subseteq\{x\,|\,\Delta^{d}f(x,\ldots,x,x)=0\}=\operatorname{Z}(f_{d})$.
+We may assume $\deg(A_{i})\geq 1$, as otherwise $A_{i}=0$ so we may discard it.
 By Corollary 3.3, $\operatorname{cor}(f_{d},P)\geq p^{-dm}$ for some polynomial $P=\sum_{i=1}^{m}c_{i}A_{i}$ of degree at most d1d-1.
-Note that $\operatorname{cor}(f,\,P-f_{<d})=\operatorname{bias}(f-(P-f_{<d}))=\operatorname{bias}(f_{d}-P)$.
+Note that $\operatorname{cor}(f,\,P+f_{<d})=\operatorname{bias}(f-(P+f_{<d}))=\operatorname{bias}(f_{d}-P)$.
 Thus,
@@
 Definition 4.1 ($\operatorname{rk}^{*}$).
-For a polynomial $f$, $\operatorname{rk}^{*}(f)$ is the smallest $r$ such that:
-  • $f=\sum_{i=1}^{r}f_{i}$ for some reducible polynomials $f_{i}=\alpha_{i}\beta_{i}$ with $\deg(f_{i})\leq\deg(f)$, and
+For a non-constant polynomial $f$, $\operatorname{rk}^{*}(f)$ is the smallest $r$ such that:
+  • $f=\sum_{i=1}^{r}f_{i}$ for some reducible polynomials $f_{i}=\alpha_{i}\beta_{i}$ with $\deg(f_{i})\leq\deg(f)$, [^17] That is, $\deg(\alpha_{i}),\deg(\beta_{i})\geq 1$ and $\deg(\alpha_{i})+\deg(\beta_{i})\leq\deg(f)$.
+Note that $\deg(\alpha_{i}),\deg(\beta_{i})<\deg(f)$.
+and
   • $\{\alpha_{i}\,|\,\deg(\alpha_{i})=1,\deg(f_{i})=\deg(f)\}$ are affinely independent, where $\deg(\alpha_{i})\leq\deg(\beta_{i})$.
-We have $\operatorname{rk}^{*}(f)=\infty$ if no such decomposition exists, for example if $\deg(f)=1$; however, as $f$ need not be homogeneous, this can happen in higher degree as well (see Remark 4.2 below).
-We note that if $f$ is homogeneous then $\operatorname{rk}^{*}(f)=\operatorname{rk}(f)$; indeed, the factors may be assumed to be linearly independent forms (see Lemma 4.6), so the affine independence condition is trivially satisfied.
-As an example for $\operatorname{rk}^{*}$, consider a polynomial of the form $f=(A+B)\cdot\ell+B$, where $\deg(A)=\deg(B)>1$ and $\deg(\ell)=1$.
-While $f$ has the decomposition $f=\ell A+(\ell+1)B$, this is not a $\operatorname{rk}^{*}$-decomposition since its degree-$1$ factors $\ell$ and $\ell+1$ are not affinely independent.
-In fact, $\operatorname{rk}^{*}(f)$ may be arbitrarily large (see Remark 4.2 next).
+If $f$ is a constant polynomial, we set $\operatorname{rk}^{*}(f)=0$.
+By convention, $\operatorname{rk}^{*}(f)=\infty$ if no such decomposition exists.
+This happens when $\deg(f)=1$, and can also happen when $\deg(f)=2$; for example, for any irreducible univariate quadratic.
+[^18] Affine independence implies at most $n$ top-degree summands having a degree-$1$ factor.
+For a univariate quadratic, this means at most $n=1$ summands.
+On the other hand, $\deg(f)\geq 3$ implies $\operatorname{rk}^{*}(f)<\infty$; in fact, $\operatorname{rk}^{*}(f)\leq n+1$.
+[^19] Indeed, let $q$ be any reducible quadratic with affine part $q_{\leq 1}=f_{\leq 1}$ (say $q(x)=(x_{1}+1)(f_{\leq 1}(x)-f_{0}x_{1})$).
+Then fqf-q has only monomials of degree at least $2$, so it has a $\operatorname{rk}^{*}$-decomposition with $\alpha_{1}=x_{1},\ldots,\alpha_{n}=x_{n}$.
+Since $\deg(q)<3\leq\deg(f)$, this gives a $\operatorname{rk}^{*}$-decomposition for $f$ of length at most $n+1$.
+(This bound cannot be improved to $n$ in general; e.g., any irreducible univariate $f$, assuming $\mathbb{F}$ is not algebraically closed, has $\operatorname{rk}^{*}(f)\geq 2=n+1$.)
+Note that if $f$ is homogeneous then $\operatorname{rk}^{*}(f)=\operatorname{rk}(f)$.
+Indeed, $\operatorname{rk}^{*}(f)\leq\operatorname{rk}(f)$ since the factors may be assumed to be linearly independent forms (see Lemma 4.6), so the affine independence condition trivially holds; conversely, $\operatorname{rk}(f)\leq\operatorname{rk}^{*}(f)$ since for each summand $\alpha_{i}\beta_{i}$, the lower-degree components of $\alpha_{i}$ and of $\beta_{i}$ contribute only monomials of degree below $\deg(f)$ to the product, and thus may all be discarded.
+Note too that $f$ of positive degree is irreducible if and only if $\operatorname{rk}^{*}(f)>1$, similarly to $\operatorname{rk}$.
+The importance of $\operatorname{rk}^{*}(f)$ stems from the fact that it is closely related to correlation with polynomials of degree $\deg(f)-2$, whereas the usual $\operatorname{rk}(f)$ only gives information about correlation with degree $\deg(f)-1$.
+As an example for $\operatorname{rk}^{*}$, consider a polynomial of the form $f=(A-B)\cdot\ell+B$, where $\deg(A)=\deg(B)>1$ and $\deg(\ell)=1$.
+While $f$ has the decomposition $f=\ell A+(-\ell+1)B$, this is not a $\operatorname{rk}^{*}$-decomposition since its degree-$1$ factors $\ell$ and $\ell+1$ are not affinely independent.
+In fact, $\operatorname{rk}^{*}(f)$ may be arbitrarily large.
+Thus, it is not surprising that $f$ may have negligible correlation with polynomials of degree at most $\deg(f)-2$ (see Remark 4.2 next).
+This is in sharp contrast to polynomials with low $\operatorname{rk}^{*}$, which do have significant correlation with polynomials of degree at most $\deg(f)-2$ (see Proposition 4.7 below).
 Remark 4.2.
-Interestingly, $\operatorname{rk}^{*}(f)$ can be more closely related to $\operatorname{bias}(f)$, or rather to analytic rank $\operatorname{AR}(f)=-\log_{p}(\operatorname{bias}(P))$, than other standard notions of rank in the literature.
+Interestingly, $\operatorname{rk}^{*}(f)$ can be more closely related to $\operatorname{bias}(f)$, or rather to analytic rank $\operatorname{AR}(f)=-\log_{|\mathbb{F}|}(\operatorname{bias}(f))$, than other standard notions of rank in the literature.
 For concreteness, consider the notion used by Green and Tao in [GT09]: $\operatorname{rank}(f)$ is the least number $k$ of polynomials $\alpha_{1},\ldots,\alpha_{k}$ with $\deg(\alpha_{i})<\deg(f)$ such that $f\equiv\Gamma(\alpha_{1},\ldots,\alpha_{k})$ for some function $\Gamma\colon\mathbb{F}^{k}\to\mathbb{F}$.
-Consider the following polynomials $f$:
-  • $f(x,y)=A(x)+\ell(y)$ with $\deg(A)>\deg(\ell)=1$.
-Clearly, $\operatorname{rank}(f)\leq\operatorname{rank}(A)+1<\infty$.
-However, $\operatorname{rk}^{*}(f)=\infty$, and moreover, $\operatorname{bias}(f)=\operatorname{bias}(A)\operatorname{bias}(\ell)=0$, so $\operatorname{AR}(f)=\infty$ as well.
-  • $f(x,y)=(A-B)\ell+B$ with $\ell=\ell(y)$ of degree $1$, and $A=A(x)$,$B=B(x)$ random of degree $d\geq 2$ on $n$ variables.
+Consider the polynomial $f(x,y)=(A-B)\ell+B$ with $\ell=\ell(y)$ of degree $1$, and $A=A(x)$, $B=B(x)$ random of constant degree $d\geq 2$ on $n$ variables.
 Clearly, $\operatorname{rank}(f)\leq 3$.
-In contrast, we claim that both $\operatorname{rk}^{*}(f)$ and $\operatorname{AR}(f)$ grow with $n$.
-    Consider a $\operatorname{rk}^{*}$-decomposition of $f$.
-By affine independence, setting all degree-$1$ factors to 0 gives a non-empty affine subspace.
-Let $U$ be the non-empty intersection of this affine subspace with the hypersurface $\ell=c$, for an appropriate $c\in\mathbb{F}$.
-Then $\operatorname{codim}(U)\leq\operatorname{rk}^{*}(f)+1$, and the restriction $f^{\prime}$ of $f$ to $U$ has $\operatorname{rank}(f^{\prime})\leq\operatorname{rk}^{*}(f)$.
-By construction, $f^{\prime}$ is a nonzero linear combination of $A$ and $B$ restricted to $U$, so $\operatorname{rank}(f^{\prime})$ grows with $n$, and thus also $\operatorname{rk}^{*}(f)$.
-    As for $\operatorname{AR}(f)$, we have $\operatorname{bias}(f)=\operatorname*{\mathbb{E}}_{x}\chi(B(x))\operatorname*{\mathbb{E}}_{y}\chi(\ell(y)(A(x)-B(x)))=\operatorname*{\mathbb{E}}_{x}\chi(B(x))\mathbf{1}_{A=B}(x)\\ =\sum_{b}\chi(b)\mathbb{P}(B=b,A=b)$, so $\operatorname{bias}(f)\approx(1/q)\sum_{b}\chi(b)\mathbb{P}(B=b)=(1/q)\operatorname{bias}(B)$, implying that $\operatorname{AR}(f)\approx n$, as claimed.
+In contrast, we will next show (slightly informally) that $\operatorname{rk}^{*}(f)$ and $\operatorname{AR}(f)$ grow with $n$, the number of $x$ variables.
+We note that a random $n$-variate polynomial over $\mathbb{F}_{q}$ of degree $d$ has exponential small $q^{-\Omega_{d}(n)}$ correlation with lower-degree polynomials [BHL12, BGY20].
+Consider a $\operatorname{rk}^{*}$-decomposition $f=\sum_{i=1}^{\operatorname{rk}^{*}(f)}\alpha_{i}\beta_{i}$.
+By affine independence, setting to 0 all degree-$1$ factors $\alpha_{i}$ coming from top-degree summands gives an affine subspace that is not empty.
+Let $U$ be the non-empty intersection of this affine subspace with the affine hyperplane $\ell=c$, for an appropriate $c\in\mathbb{F}$.
+Note that $\operatorname{codim}(U)\leq\operatorname{rk}^{*}(f)+1$.
+Observe that the restriction $f^{\prime}$ of $f$ to $U$ has (in the sense of Green-Tao) $\operatorname{rank}(f^{\prime})\leq 2\operatorname{rk}^{*}(f)$, since $\deg(f^{\prime})=d$ (with high probability) and it inherits a decomposition in terms of polynomials of degree at most $\deg(f)-2=d-1$.
+By construction, $f^{\prime}$ is a nonzero linear combination of $A$ and $B$ restricted to $U$; using our bound on $\operatorname{codim}(U)$ and the randomness of $A$ and $B$, $\operatorname{rank}(f^{\prime})$ grows with $n$, and thus also $\operatorname{rk}^{*}(f)$.
+We will next see that $\operatorname{cor}(f,P)\leq|\mathbb{F}|^{-\Omega(n)}$ for any $P\in\operatorname{Poly}_{\deg(f)-2}$. Note that
+$\operatorname{cor}(f,P)=|\operatorname*{\mathbb{E}}_{y}\operatorname*{\mathbb{E}}_{x}\chi(f(x,y)-P(x,y))|\leq\max_{y}|\operatorname*{\mathbb{E}}_{x}\chi(f(x,y)-P(x,y))|=\max_{y}\operatorname{cor}(f_{y},P_{y})$
+where $f_{y}(x)=f(x,y)$ and $P_{y}(x)=P(x,y)$.
+For every fixed $y$, we have by construction that $f_{y}$ is a nontrivial linear combination of two random polynomials of degree $d$, and thus is a random polynomial of degree $d$ as well; therefore, since $\deg(P_{y})\leq\deg(P)\leq d-1$, we have that $\operatorname{cor}(f_{y},P_{y})\leq|\mathbb{F}|^{-\Omega(n)}$.
+By the above, the same follows for $\operatorname{cor}(f,P)$.
+[^20] An alternative proof that $\operatorname{rk}^{*}(f)$ grows with $n$ follows immediately by combining this correlation upper bound together with Proposition 4.7 below.
+Finally, note that this correlation upper bound in particular applies to $\operatorname{bias}(f)=\operatorname{cor}(f,0)$, so $\operatorname{AR}(f)\geq\Omega(n)$, similarly to $\operatorname{rk}^{*}(f)$.
 4.1 Properties of $\operatorname{rk}^{*}$
@@
 Observation 4.3 (subadditivity).
-$\operatorname{rk}^{*}(g+h)\leq\operatorname{rk}^{*}(g)+\operatorname{rk}^{*}(h)$ if $\deg(h)\neq\deg(g)$.
+$\operatorname{rk}^{*}(g+h)\leq\operatorname{rk}^{*}(g)+\operatorname{rk}^{*}(h)$ for positive $\deg(h)\neq\deg(g)$.
 A basic fact about rank is that it is not inherited by partial derivatives: even if a form has low rank, a partial derivative might not.
 Consider, say, reducible forms; for example, we have $\operatorname{rk}(x\cdot S)\leq 1$ for any form $S=S(y)$, yet the partial derivative with respect to $x$ is $S(y)$, which may have arbitrarily high rank.
-That being said, we show that, perhaps surprisingly, adding to a polynomial a partial derivative does not increase its $\operatorname{rk}^{*}$.
+That being said, we show that, perhaps surprisingly, adding to a polynomial a partial derivative does not increase its $\operatorname{rk}^{*}$, up to a polynomial of lower degree.
+[^21] Alternatively, one could omit the lower degree polynomial $\gamma$ from Lemma 4.4, and from Proposition 4.5 below, by redefining $\operatorname{rk}^{*}(f)$ to ignore the components of $f$ of degree at most $\deg(f)-2$.
 Lemma 4.4 (derivation invariance).
-For any form $g$ and any direction vector $c$,
-$\operatorname{rk}^{*}(g+\partial_{c}g)\leq 2\operatorname{rk}(g).$
+For any $g\in\operatorname{Form}_{k}$ with $k\geq 2$, and any direction vector $c$,
+$\operatorname{rk}^{*}(g+\partial_{c}g+\gamma)\leq\operatorname{rk}(g)$
+for some $\gamma\in\operatorname{Poly}_{k-2}$.
+For example, for the univariate $g(x)=x^{2}\in\operatorname{Form}_{2}$, we have $\partial_{c}g=2cx$, and it is easy to see that there is a constant $\gamma\in\mathbb{F}$ such that $g+\partial_{c}g+\gamma$ is reducible (is of $\operatorname{rk}^{*}$ 11); indeed, take $\gamma=c^{2}$.
+Lemma 4.4 can be thought of as a generalization of such identities.
 Combining the above two properties, we show that $\operatorname{rk}^{*}$ satisfies a certain approximate “chain rule”, which would complement a chain rule for bias that we prove later in Section 6.
 Proposition 4.5 (chain rule for $\operatorname{rk}^{*}$).
-For any polynomial $g+h$ with top-degree component $g$, and any direction vector $c$,
-$\operatorname{rk}^{*}(g+h)\leq 2\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g).$
-We now prove the above properties of $\operatorname{rk}^{*}$.
+For any $g\in\operatorname{Form}_{k}$ with $k\geq 2$ and $h\in\operatorname{Poly}_{<k}$, and any direction vector $c$,
+$\operatorname{rk}^{*}(g+h+\gamma)\leq\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g)$
+for some $\gamma\in\operatorname{Poly}_{k-2}$.
+We now prove the above properties of $\operatorname{rk}^{*}$. We start with subadditivity.
 Proof of Observation 4.3.
-Put $r_{1}=\operatorname{rk}^{*}(g)$, $r_{2}=\operatorname{rk}^{*}(h)$.
-Let $g=\sum_{i=1}^{r_{1}}g_{i}$ and $h=\sum_{i=1}^{r_{2}}h_{i}$, with $g_{i}$ and $h_{i}$ reducible polynomials with $\deg(g_{i})\leq\deg(g)$ and $\deg(h_{i})\leq\deg(h)$, be decompositions witnessing $\operatorname{rk}^{*}(g)$ and $\operatorname{rk}^{*}(h)$, respectively.
+Put $r_{1}=\operatorname{rk}^{*}(g)$, $r_{2}=\operatorname{rk}^{*}(h)$, and assume both are finite as otherwise there is nothing to prove.
+Let $g=\sum_{i=1}^{r_{1}}g_{i}$ and $h=\sum_{i=1}^{r_{2}}h_{i}$, with $g_{i}$ and $h_{i}$ reducible polynomials with $\deg(g_{i})\leq\deg(g)$ and $\deg(h_{i})\leq\deg(h)$, be decompositions witnessing $\operatorname{rk}^{*}(g)$ and $\operatorname{rk}^{*}(h)$, respectively (recall that $g$ and $h$ are not constant).
 We have $g+h=\sum_{i=1}^{r_{1}}g_{i}+\sum_{i=1}^{r_{2}}h_{i}$.
 Assume without loss of generality that $\deg(h)<\deg(g)$.
-Since $\deg(h_{i})\leq\deg(h)$, observe that the affine independence condition for $\operatorname{rk}^{*}(g+h)$ applies only for the $g_{i}$, and is therefore satisfied by the definition of $\operatorname{rk}^{*}(g)$.
+Since $\deg(h_{i})\leq\deg(h)$, observe that the affine independence condition for $\operatorname{rk}^{*}(g+h)$ applies only for the $g_{i}$ (of degree $\deg(g)$ with a degree-$1$ factor), and is therefore satisfied by the definition of $\operatorname{rk}^{*}(g)$.
 Thus, $\operatorname{rk}^{*}(g+h)\leq r_{1}+r_{2}=\operatorname{rk}^{*}(g)+\operatorname{rk}^{*}(h)$, as needed.
 ∎
+We will use the following easy observation.
 Lemma 4.6.
@@
 Observe that $g$ has the decomposition $g=\sum_{i=1}^{r-1}\alpha_{i}(\beta_{i}+c_{i}\beta_{r})$.
-Each summand is a reducible form, since $\alpha_{i}$ is homogeneous as well as $\beta_{i}+c_{i}\beta_{r}$; indeed, $\deg(\beta_{i})=k-\deg(\alpha_{i})=k-\deg(\alpha_{r})=\deg(\beta_{r})$ whenever $c_{i}\neq 0$.
+Each summand is a reducible form, since $\alpha_{i}$ is homogeneous as well as $\beta_{i}+c_{i}\beta_{r}$; indeed, $\deg(\beta_{i})=\deg(g)-\deg(\alpha_{i})=\deg(g)-\deg(\alpha_{r})=\deg(\beta_{r})$ whenever $c_{i}\neq 0$.
 Thus, $r=\operatorname{rk}(g)\leq r-1$, a contradiction.
@@
 Proof of Lemma 4.4.
-Put $r=\operatorname{rk}(g)$ and $k=\deg(g)$, and write $g=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ with forms $\alpha_{i},\beta_{i}$ satisfying $\deg(\alpha_{i})\leq\deg(\beta_{i})<k$ and $\deg(\alpha_{i}\beta_{i})=k$.
+Put $r=\operatorname{rk}(g)$ and write $g=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ with forms $\alpha_{i},\beta_{i}$ satisfying $1\leq\deg(\alpha_{i})\leq\deg(\beta_{i})<k$ and $\deg(\alpha_{i}\beta_{i})=k$.
 By the product rule, $\partial g_{c}=\sum_{i=1}^{r}\alpha_{i}(\partial_{c}\beta_{i})+(\partial_{c}\alpha_{i})\beta_{i}$. Observe that we have the following fortunate identity:
 $g+\partial_{c}g=\sum_{i=1}^{r}(\alpha_{i}+\partial_{c}\alpha_{i})(\beta_{i}+\partial_{c}\beta_{i})-\sum_{i=1}^{r}(\partial_{c}\alpha_{i})(\partial_{c}\beta_{i}).$
-Now, since $\deg(\alpha_{i}+\partial_{c}\alpha_{i})=\deg(\alpha_{i})\leq\deg(\beta_{i})=\deg(\beta_{i}+\partial_{c}\beta_{i})$, the linear components of the factors $\alpha_{i}+\partial_{c}\alpha_{i}$ are precisely the linear factors $\alpha_{i}$ in the above decomposition for $g$, which by Lemma 4.6 are linearly independent.
-As for the remaining terms, note that $\deg((\partial_{c}\alpha_{i})(\partial_{c}\beta_{i}))\leq\deg(\alpha_{i}\beta_{i})-2=k-2<\deg(g+\partial_{c}g)$, so $\operatorname{rk}^{*}(g+\partial_{c}g)$ has no requirement on their (linear) factors.
-We deduce that the above (non-homogeneous) decomposition is a $\operatorname{rk}^{*}$-decomposition, so $\operatorname{rk}^{*}(g+\partial_{c}g)\leq 2r=2\operatorname{rk}(g)$, as needed.
+Put $\alpha_{i}^{\prime}=\alpha_{i}+\partial_{c}\alpha_{i}$ and $\beta_{i}^{\prime}=\beta_{i}+\partial_{c}\beta_{i}$, and note that $\deg(\alpha_{i}^{\prime})=\deg(\alpha_{i})\leq\deg(\beta_{i})=\deg(\beta_{i}^{\prime})$.
+Thus, the linear components of those factors $\alpha^{\prime}_{i}$ with $\deg(\alpha^{\prime}_{i})=1$ are precisely the $\alpha_{i}$ with $\deg(\alpha_{i})=1$, which by Lemma 4.6 are linearly independent.
+As for the remaining terms, note that $\deg((\partial_{c}\alpha_{i})(\partial_{c}\beta_{i}))\leq\deg(\alpha_{i}\beta_{i})-2=k-2$, so their sum $\gamma$ has $\deg(\gamma)\leq k-2$.
+Thus, we have a (non-homogeneous) $\operatorname{rk}^{*}$-decomposition of $g+\partial_{c}g+\gamma$, and thus $\operatorname{rk}^{*}(g+\partial_{c}g+\gamma)\leq r=\operatorname{rk}(g)$, as needed.
+∎
 We are now ready to prove our chain rule for $\operatorname{rk}^{*}$.
 Proof of Proposition 4.5.
-We have
-$\operatorname{rk}^{*}(g+h)$ $=\operatorname{rk}^{*}\big((g+\partial_{c}g)+(h-\partial_{c}g)\big)$
-$\leq\operatorname{rk}^{*}(g+\partial_{c}g)+\operatorname{rk}^{*}(h-\partial_{c}g)\leq 2\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g)$
-where the first inequality uses the subadditivity of $\operatorname{rk}^{*}$ in Observation 4.3, and the second inequality uses Lemma 4.4. ∎
+Apply Lemma 4.4 on $g\in\operatorname{Form}_{k}$ and $c$ to obtain $\gamma\in\operatorname{Poly}_{k-2}$ with $\operatorname{rk}^{*}(g+\partial_{c}g+\gamma)\leq\operatorname{rk}(g)$.
+If $h-\partial_{c}g$ is a constant $a$, then we are done since $\gamma-a\in\operatorname{Poly}_{k-2}$ and
+$\operatorname{rk}^{*}(g+h+(\gamma-a))=\operatorname{rk}^{*}(g+\partial_{c}g+\gamma)\leq\operatorname{rk}(g)=\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g).$
+Otherwise,
+$\operatorname{rk}^{*}(g+h+\gamma)$ $=\operatorname{rk}^{*}\big((g+\partial_{c}g+\gamma)+(h-\partial_{c}g)\big)$
+$\leq\operatorname{rk}^{*}(g+\partial_{c}g+\gamma)+\operatorname{rk}^{*}(h-\partial_{c}g)\leq\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g)$
+where the first inequality uses the subadditivity of $\operatorname{rk}^{*}$ in Observation 4.3, as the two polynomials have different positive degrees: $k$ and at most k1k-1. ∎
 4.2 Lower-degree correlation
@@
 In this section we show that if a polynomial $f$ has a short $\operatorname{rk}^{*}$-decomposition, then it has a significant correlation with a polynomial of degree at most $\deg(f)-2$.
+[^22] Assuming $\deg(f)\geq 3$.
+If $\deg(f)=2$ we only get correlation with degree $\deg(f)-1$, naturally.
 This relies on our correlation lemma in Lemma 3.1, but here we only assume small $\operatorname{rk}^{*}$.
@@
 $\operatorname{cor}_{<k-1}(f)\geq q^{-2\operatorname{rk}^{*}(f)}.$
-In order to convert a short $\operatorname{rk}^{*}$-decomposition to significant correlation, via Lemma 3.1, we use the following lemma which gives an essentially optimal lower bound on the joint zero probability $\mathbb{P}(\operatorname{Z}(\mathbf{A}))$.
+For completeness, if $\operatorname{rk}^{*}(f)=\infty$ then the right hand side should be interpreted as 0.
+In order to convert a short $\operatorname{rk}^{*}$-decomposition to significant correlation, via Lemma 3.1, we use the following lemma which puts a shift of $f$ in an ideal generated by few polynomials of degree at most $\deg(f)-2$ that, moreover, frequently vanish together.
 Lemma 4.8 (perturbation lemma).
@@
 Proof.
-Put $r=\operatorname{rk}^{*}(P)$ and $\mathbb{F}=\mathbb{F}_{q}$.
+Put $r=\operatorname{rk}^{*}(f)$ and $\mathbb{F}=\mathbb{F}_{q}$.
+Assume $r<\infty$, as otherwise we are done by taking $\mathbf{A}=(x_{1},\ldots,x_{n})$ and $c=f(0)$.
+Assume $f$ is non-constant, as otherwise we are done by taking $m=0$ and $c=f\in\mathbb{F}_{q}$.
 Write $f=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ with $1\leq\deg(\alpha_{i})\leq\deg(\beta_{i})\leq k-1$ and $\deg(\alpha_{i}\beta_{i})\leq k$.
-Assume without loss of generality that all polynomials $\alpha_{i}$ satisfying both $\deg(\alpha_{i})=1$ and $\deg(\alpha_{i}\beta_{i})=k$ are precisely $\alpha_{t+1},\ldots,\alpha_{r}$ for some $t\leq r$.
-Observe that for every $i\leq t$ we necessarily have $\deg(\beta_{i})=\deg(\alpha_{i}\beta_{i})-\deg(\alpha_{i})\leq k-2$.
-Put $\mathbf{A}^{1}=(\alpha_{1},\ldots,\alpha_{t})$, $\mathbf{A}^{2}=(\alpha_{t+1},\ldots,\alpha_{r})$, and $\mathbf{B}^{1}=(\beta_{1},\ldots,\beta_{t})$. Since
+Assume without loss of generality that the polynomials $\alpha_{i}$ satisfying both $\deg(\alpha_{i})=1$ and $\deg(\alpha_{i}\beta_{i})=k$ are precisely $\alpha_{t+1},\ldots,\alpha_{r}$ for some $t\leq r$.
+Observe that for every $i\leq t$ we necessarily have $\deg(\beta_{i})=\deg(\alpha_{i}\beta_{i})-\deg(\alpha_{i})\leq k-2$; indeed, we have that $\deg(\alpha_{i}\beta_{i})\leq k-1$ or $\deg(\alpha_{i})\geq 2$.
+Let $\mathbf{A}^{1}=(\alpha_{1},\ldots,\alpha_{t})$, $\mathbf{A}^{2}=(\alpha_{t+1},\ldots,\alpha_{r})$, and $\mathbf{B}^{1}=(\beta_{1},\ldots,\beta_{t})$, and $m=r+t$.
+We claim that there are $y^{1},z^{1}\in\mathbb{F}^{t}$ such that the map $(\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ satisfies
+$\mathbb{P}\big((\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})=(y^{1},0,z^{1})\big)\geq q^{-m}.$ (4)
+If $t=r$ (for example, when $\deg(f)<k$), that is, $\mathbf{A}^{2}$ is empty, then (4) follows by the pigeonhole principle, so henceforth assume otherwise.
+By the definition of $\operatorname{rk}^{*}(f)$, the linear components of $\alpha_{t+1},\ldots,\alpha_{r}$ are linearly independent, which implies that the affine map $\mathbf{A}^{2}\colon\mathbb{F}^{n}\to\mathbb{F}^{r-t}$ is perfectly equidistributed; indeed, the associated linear map is surjective, so its fibers are all cosets of the kernel and thus of the same size.
+In particular, $\mathbb{P}(\mathbf{A}^{2}=0)=q^{-(r-t)}$ (>0>0).
+Since
 $\sum_{y,z\in\mathbb{F}^{t}}\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y,z)\,|\,\mathbf{A}^{2}=0\big)=1,$
 there exist $y^{1},z^{1}\in\mathbb{F}^{t}$ such that
-$\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y^{1},z^{1})\,|\,\mathbf{A}^{2}=0\big)\geq q^{-2t}.$ (4)
-Put $y=(y^{1},y^{2})\in\mathbb{F}^{r}$ with $y^{2}=0\in\mathbb{F}^{r-t}$. We decompose $f$ as follows (recall $y_{i}=0$ for $i\geq t$):
+$\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y^{1},z^{1})\,|\,\mathbf{A}^{2}=0\big)\geq q^{-2t}.$
+This proves (4):
+$\mathbb{P}\big((\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})=(y^{1},0,z^{1})\big)$ $=\mathbb{P}(\mathbf{A}^{2}=0)\cdot\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y^{1},z^{1})\,|\,\mathbf{A}^{2}=0\big)$
+$\geq q^{-(r-t)}q^{-2t}=q^{-m}.$
+Now, put $y=(y^{1},y^{2})\in\mathbb{F}^{r}$ with $y^{2}=0\in\mathbb{F}^{r-t}$. Decompose $f$ as follows (recall $y_{i}=0$ for $i>t$):
 $f=\sum_{i=1}^{r}(\alpha_{i}-y_{i})\beta_{i}+\sum_{i=1}^{t}y_{i}(\beta_{i}-z_{i})+\sum_{i=1}^{t}y_{i}z_{i}.$
@@
 Then $f-c\in\langle A_{1},...,A_{m}\rangle$ with $m=r+t\leq 2r$.
-Moreover, for $i\leq k$ we have $\deg(A_{r+i})=\deg(\beta_{i})\leq k-2$, and for $i\leq r$ we have $\deg(A_{i})=\deg(\alpha_{i})\leq\lfloor k/2\rfloor\leq k-2$, using our assumption $k\geq 3$.
-Finally, by the definition of $\operatorname{rk}^{*}(f)$, the linear components of $\alpha_{t+1},\ldots,\alpha_{r}$ are linearly independent.
-Thus, the affine map $\mathbf{A}^{2}\colon\mathbb{F}^{n}\to\mathbb{F}^{r-t}$ is perfectly equidistributed, and in particular, $\mathbb{P}(\mathbf{A}^{2}=0)\geq q^{-(r-t)}$.
-Therefore,
-$\mathbb{P}(\forall i\colon A_{i}=0)$ $=\mathbb{P}((\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})=(y^{1},y^{2},z^{1}))$
-$=\mathbb{P}(\mathbf{A}^{2}=0)\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y^{1},z^{1})\,|\,\mathbf{A}^{2}=0\big)\geq q^{-(r-t)}q^{-2t}=q^{-m},$
-where the inequality combines the equidistribution of $\mathbf{A}^{2}$ together with (4). This completes the proof. ∎
+Moreover, for $i\leq t$ we have $\deg(A_{r+i})=\deg(\beta_{i})$, and $1\leq\deg(\beta_{i})\leq k-2$, while for $i\leq r$ we have $\deg(A_{i})=\deg(\alpha_{i})$, and $1\leq\deg(\alpha_{i})\leq\lfloor k/2\rfloor\leq k-2$ using our assumption $k\geq 3$.
+Observe that
+$\mathbb{P}(\forall i\colon A_{i}=0)=\mathbb{P}\big((\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})=(y^{1},y^{2},z^{1})\big)\geq q^{-m},$
+where the inequality is (4). This completes the proof. ∎
 Proof of Proposition 4.7.
@@
 […]
-Since $\sum_{i=1}^{m}c_{i}A_{i}\in\operatorname{Poly}_{k-2}$, we deduce that
+Since $\sum_{i=1}^{m}c_{i}A_{i}\in\operatorname{Poly}_{k-2}$, we deduce that $f$ is significantly correlated with a polynomial of degree at most k2k-2:
 $\operatorname{cor}_{<k-1}(f)\geq\max_{c\in\mathbb{F}^{m}}\operatorname{cor}\Big(f,\,\sum_{i=1}^{m}c_{i}A_{i}\Big)\geq q^{-2\operatorname{rk}^{*}(f)}.\qed$
@@
 For any form $f(x)=\sum_{I}c_{I}\prod_{j=1}^{k}x_{I_{j}}$, with each $I=(I_{1},\ldots,I_{k})$ a sorted $k$-tuple of variables and $c_{I}$ a scalar,
-$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})=\sum_{I}c_{I}\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{I_{j}}^{(\varphi(j))}.$
+$\Delta^{d}_{x^{(1)},\ldots,x^{(d)}}f(x^{(d+1)})=\sum_{I}c_{I}\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{I_{j}}^{(\varphi(j))}.$
 Proof.
 By linearity we may assume that $f$ is a single monomial, $f=\prod_{j=1}^{k}x_{i_{j}}$, so it suffices to prove
-$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})=\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{i_{j}}^{(\varphi(j))}.$
+$\Delta^{d}_{x^{(1)},\ldots,x^{(d)}}f(x^{(d+1)})=\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{i_{j}}^{(\varphi(j))}.$
 We proceed by induction on $d\geq 0$.
 The base case $d=0$ holds since the only function $\varphi\colon[k]\to\{1\}$ is $\varphi=1$, so the right hand side is $\prod_{j=1}^{k}x_{i_{j}}^{(1)}=f(x^{(1)})=\Delta^{0}_{\mathbf{x}}f(x^{(1)})$.
-For the induction step, we have
-$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$ $=\Delta_{x^{(d)}}\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)})=\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)}+x^{(d)})-\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)})$
+For the induction step, abbreviate $\mathbf{x}=(x^{(1)},\ldots,x^{(d-1)})$, and observe that
+$\Delta^{d}_{x^{(1)},\ldots,x^{(d)}}f(x^{(d+1)})$ $=\Delta_{x^{(d)}}\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)})=\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)}+x^{(d)})-\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)})$
 $=\sum_{\varphi\colon[k]\twoheadrightarrow[d-1]}\,\prod_{j\colon\varphi(j)\neq d}x_{i_{j}}^{(\varphi(j))}\Big(\prod_{j\colon\varphi(j)=d}(x_{i_{j}}^{(d+1)}+x_{i_{j}}^{(d)})-\prod_{j\colon\varphi(j)=d}x_{i_{j}}^{(d+1)}\Big)$
@@
 $=\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{i_{j}}^{(\varphi(j))},$
-where the third equality uses the induction hypothesis, and the last equality uses the natural bijection between semi-surjections $\varphi^{\prime}\colon[k]\twoheadrightarrow[d]$ and pairs of semi-surjections $(\varphi,\tau)$ where $\varphi\colon[k]\twoheadrightarrow[d-1]$ and $\tau\colon\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle\!-\!1$}}(d)\to\{d,d+1\}$ with $d\in\operatorname{Im}(\tau)$.
+where the third equality uses the induction hypothesis twice, and the last equality uses the natural bijection between semi-surjections $\varphi^{\prime}\colon[k]\twoheadrightarrow[d]$ and pairs $(\varphi,\tau)$ where $\varphi\colon[k]\twoheadrightarrow[d-1]$ and $\tau\colon\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle\!-\!1$}}(d)\to\{d,d+1\}$ with $d\in\operatorname{Im}(\tau)$.
 This completes the induction step and the proof.
@@
 For any monomial $g(x)=\prod_{j=1}^{k}x_{i_{j}}$,
-$\overline{g}(x^{(1)},\ldots,x^{(k)})=\sum_{\pi\in S_{k}}\,\prod_{j=1}^{k}x_{i_{j}}^{(\pi(j))}\,.$
+$\overline{g}(x^{(1)},\ldots,x^{(k)})=\sum_{\sigma\in S_{k}}\,\prod_{j=1}^{k}x_{i_{j}}^{(\sigma(j))}\,.$
 Fact 5.7.
@@
 Proof of Lemma 5.2.
-It suffices to show $\partial_{x_{n}^{(k)}}\overline{g}(\mathbf{x})=\overline{\partial_{x_{n}}g}(\mathbf{x})$ for $g(x)=\prod_{j=1}^{k}x_{i_{j}}$ ($i_{1}\leq\cdots\leq i_{k}$) a single monomial, and $\mathbf{x}=(x^{(1)},\ldots,x^{(k)})$, since the result would follow by symmetry and linearity.
+It suffices to show $\partial_{x_{n}^{(k)}}\overline{g}(\mathbf{x})=\overline{\partial_{x_{n}}g}(\mathbf{x})$ for $g(x)=\prod_{j=1}^{k}x_{i_{j}}$ ($i_{1}\leq\cdots\leq i_{k}$) a single monomial, and $\mathbf{x}=(x^{(1)},\ldots,x^{(k-1)})$, since the result would follow by symmetry and linearity (the polarization $\overline{g}$ is symmetric by Fact 5.6; directional derivatives are linear in $c$).
 On the one hand, [^23] For example, $\partial_{x_{2}}(x_{1}x_{2}^{2})=x_{1}x_{2}+x_{1}x_{2}$.
@@
 $\overline{g}(x^{(1)},\ldots,x^{(k)})=\sum_{\sigma\in S_{k}}\,\prod_{j=1}^{k}x_{i_{j}}^{(\sigma(j))}$
-is multilinear, so its only monomials containing $x^{(k)}_{n}$ are those where an occurrence of $x_{n}$ is replaced by $x_{n}^{(k)}$; therefore,
-$\partial_{x^{(k)}_{n}}\overline{g}(x^{(1)},\ldots,x^{(k)})=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\sum_{\begin{subarray}{c}\sigma\in S_{k}\\ \pi(t)=k\end{subarray}}\,\prod_{\begin{subarray}{c}j=1\\ j\neq t\end{subarray}}^{k}x_{i_{j}}^{(\sigma(j))}=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\sum_{\begin{subarray}{c}\sigma\in S_{k}\\ \pi(t)=k\end{subarray}}\,\prod_{j=1}^{k-1}x_{i_{j}}^{(\sigma(j))}\;.$ (6)
+is multilinear, so its only monomials containing $x^{(k)}_{n}$ are those where an occurrence of $x_{n}$ is replaced by $x_{n}^{(k)}$ (that is, $i_{t}=n$ and $\sigma(t)=k$); therefore,
+$\partial_{x^{(k)}_{n}}\overline{g}(x^{(1)},\ldots,x^{(k-1)})=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\sum_{\begin{subarray}{c}\sigma\in S_{k}\\ \sigma(t)=k\end{subarray}}\,\prod_{\begin{subarray}{c}j=1\\ j\neq t\end{subarray}}^{k}x_{i_{j}}^{(\sigma(j))}=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\sum_{\begin{subarray}{c}\sigma\in S_{k}\\ \sigma(t)=k\end{subarray}}\,\prod_{j=1}^{k-1}x_{i_{j}}^{(\sigma(j))}\;.$ (6)
 A quick inspection shows that (5) and (6) are equal, as needed.
@@
 Proof of Proposition 5.1.
-By linearity, we may assume that $f$ is a single monomial, $f=\prod_{j=1}^{d+1}x_{i_{j}}$, of degree $d+1$.
-For a function $\varphi\colon[k]\to[d+1]$, we abbreviate $f^{(\varphi)}=\prod_{j=1}^{d+1}x_{i_{j}}^{(\varphi(j))}$.
-By Lemma 5.5,
+By linearity, we may assume that we are given a single monomial, $f=\prod_{j=1}^{d+1}x_{i_{j}}$, of degree $d+1$. We need to show, for $\mathbf{x}=(x^{(1)},\ldots,x^{(d)})$, that
+$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})=\overline{f}\Big(\mathbf{x},\,x^{(d+1)}+\frac{1}{2}\sum_{t=1}^{d}x^{(t)}\Big).$
+For a function $\varphi\colon[d+1]\to[d+1]$, we abbreviate $f^{(\varphi)}=\prod_{j=1}^{d+1}x_{i_{j}}^{(\varphi(j))}$.
+Observe that every semi-surjection $\varphi\colon[d+1]\twoheadrightarrow[d]$ is either a permutation or takes exactly one value twice.
+Therefore, by Lemma 5.5,
 $\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$ $=\sum_{\varphi\colon[d+1]\twoheadrightarrow[d]}\,f^{(\varphi)}$
@@
 where $S_{d}^{t}$ denotes the set of all surjections $\varphi\colon[d+1]\to[d]$ satisfying $|\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle\!-\!1$}}(t)|=2$.
-We claim that for any $t\in[d]$, there is a 22-to-11 map from $S_{d+1}$ to $S_{d}^{t}$.
-Indeed, we map each permutation $\sigma\in S_{d+1}$ to the function $\sigma_{t}\in S_{d}^{t}$ given by
+For $t\in[d]$ and a permutation $\sigma\in S_{d+1}$, consider the function $\sigma_{t}\in S_{d}^{t}$ given by
 $\sigma_{t}(x):=\begin{cases}t&\text{if }\sigma(x)=d+1\\ \sigma(x)&\text{else}\end{cases}\;.$
-Put differently, $\sigma_{t}(x)=\sigma_{\sigma^{\raisebox{0.43054pt}{$\scriptscriptstyle\!-\!1$}}(d+1)\to t}$, where for any function $h$ we write $h_{x_{0}\to y_{0}}$ for
-$h_{x_{0}\to y_{0}}(x)=\begin{cases}y_{0}&\text{if }x=x_{0}\\ h(x)&\text{else}\end{cases}\;.$
-To see that this map is 22-to-11, let $\varphi\in S_{d}^{t}$, and consider the two functions $\varphi_{a\mapsto d+1}$ and $\varphi_{b\mapsto d+1}$.
-That both functions are in $S_{d+1}$ follows from our assumption on $\varphi$.
-A moment’s thought reveals that both—and none others—are mapped to $\varphi$, that is, $\sigma_{\sigma^{\raisebox{0.43054pt}{$\scriptscriptstyle\!-\!1$}}(d+1)\to t}=\varphi$ if and only if $\sigma=\varphi_{a\mapsto d+1}$ or $\sigma=\varphi_{b\mapsto d+1}$, proving our claim.
-Going back to $\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$, we thus have
+We claim that the map $\sigma\mapsto\sigma_{t}$, from $S_{d+1}$ to $S_{d}^{t}$, is 22-to-11. Indeed, for any $\varphi\in S_{d}^{t}$, write $\varphi^{-1}(t)=\{a,b\}$ and consider the two permutations
+$\varphi_{a\mapsto d+1}(x)=\begin{cases}d+1&\text{if }x=a\\ \varphi(x)&\text{else}\end{cases}\quad\text{ and }\quad\varphi_{b\mapsto d+1}(x)=\begin{cases}d+1&\text{if }x=b\\ \varphi(x)&\text{else}\end{cases}.$
+A moment’s thought reveals that $\sigma\in S_{d+1}$ satisfies $\sigma_{t}=\varphi$ if and only if $\sigma\in\{\varphi_{a\mapsto d+1},\varphi_{b\mapsto d+1}\}$.
+Going back to $\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$, and using the fact that $\operatorname{ch}(\mathbb{F})\neq 2$, we thus have
+$\sum_{\varphi\in S_{d}^{t}}\,f^{(\varphi)}=\frac{1}{2}\sum_{\sigma\in S_{d+1}}f^{(\sigma_{t})}.$
+Therefore,
 $\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$ $=\sum_{\sigma\in S_{d+1}}\,f^{(\sigma)}\,+\,\sum_{t=1}^{d}\,\frac{1}{2}\sum_{\sigma\in S_{d+1}}f^{(\sigma_{t})}$
@@
 $=\overline{f}\Big(\mathbf{x},x^{(d+1)}+\frac{1}{2}\sum_{t=1}^{d}x^{(t)}\Big).$
-using the multilinearity of $\overline{f}$ in the last equality. This completes the proof. ∎
+using Fact 5.6 in the second equality, and the multilinearity of $\overline{f}$ in the last equality. This completes the proof. ∎
 6 Proof of Theorem 1
@@
 Remark 6.1.
-While the bias of a polynomial is closely related to the bias of its top-degree component, it generally has little to do with the bias of its lower-degree components.
+While the bias of a polynomial is closely related to the bias of its top-degree component, it can have little to do with the bias of its lower-degree components.
 For example, consider the univariate polynomial $x^{2}+x$, over $\mathbb{F}_{p}$ with $p>2$.
 We have
-$\operatorname{bias}(x^{2}+x)\cdot p=\Big|\sum_{x\in\mathbb{F}_{p}}\omega_{p}^{x^{2}+x}\Big|=\Big|\omega_{p}^{-1/4}\sum_{x\in\mathbb{F}_{p}}\omega_{p}^{(x+1/2)^{2}}\Big|=\Big|\sum_{y\in\mathbb{F}_{p}}\omega_{p}^{y^{2}}\Big|=\sqrt{p},$
+$\operatorname{bias}(x^{2}+x)\cdot p=\Big|\sum_{x\in\mathbb{F}_{p}}e(x^{2}+x)\Big|=\Big|e(-1/4)\sum_{x\in\mathbb{F}_{p}}e((x+1/2)^{2})\Big|=\Big|\sum_{y\in\mathbb{F}_{p}}e(y^{2})\Big|=\sqrt{p},$
 where the last step uses the formula for a basic quadratic Gauss sum.
-Thus, $\operatorname{bias}(x^{2}+x)=\operatorname{bias}(x^{2})=p^{-1/2}$, whereas for the lower degree component of $x^{2}+x$ we have $\operatorname{bias}(x)=0$.
+Thus, $\operatorname{bias}(x^{2}+x)=\operatorname{bias}(x^{2})=p^{-1/2}$.
+However, for the lower degree component of $x^{2}+x$ we have $\operatorname{bias}(x)=0$.
 Remark 6.2.
@@
 Proof.
-Observe that for any fixed $y\in E^{c}$, $A(x,y)+B(y)$ is polynomial in $x$ of degree exactly $1$, and thus takes any value in the field with the same probability. Furthermore observe that
-$\operatorname{bias}(A)=\operatorname*{\mathbb{E}}(\chi(A))=\operatorname*{\mathbb{E}}_{y}\operatorname*{\mathbb{E}}_{x}(\chi(A(x,y)))=\operatorname*{\mathbb{E}}_{y}\mathbf{1}_{E}(y)=\mathbb{P}(E).$
+Observe that for any fixed $y\in E^{c}$, $A(x,y)+B(y)$ has degree exactly $1$ (in $x$), and thus takes any value in the field with the same probability. Furthermore observe that
+$\operatorname{\mathfrak{bias}}(A)=\operatorname*{\mathbb{E}}(\chi(A))=\operatorname*{\mathbb{E}}_{y}\operatorname*{\mathbb{E}}_{x}(\chi(A(x,y)))=\operatorname*{\mathbb{E}}_{y}\mathbf{1}_{E}(y)=\mathbb{P}(E).$
 Therefore, by the law of total expectation,
-$\operatorname{bias}(A+B)$ $=\operatorname*{\mathbb{E}}(\chi(A+B))=\mathbb{P}(E)\operatorname*{\mathbb{E}}(\chi(A+B)\,|\,E)+\mathbb{P}(E^{c})\operatorname*{\mathbb{E}}(\chi(A+B)\,|\,E^{c})$
-$=\operatorname{bias}(A)\operatorname{bias}(B\,|\,E)+\mathbb{P}(E^{c})\cdot 0.\qed$
-Our main result in this subsection bounds the bias of the $d$-times iterated derivative of $f\in\operatorname{Poly}_{d+1}$ in terms of the components of $f$.
+$\operatorname{\mathfrak{bias}}(A+B)$ $=\operatorname*{\mathbb{E}}(\chi(A+B))=\mathbb{P}(E)\operatorname*{\mathbb{E}}(\chi(A+B)\,|\,E)+\mathbb{P}(E^{c})\operatorname*{\mathbb{E}}(\chi(A+B)\,|\,E^{c})$
+$=\operatorname{\mathfrak{bias}}(A)\operatorname{\mathfrak{bias}}(B\,|\,E)+\mathbb{P}(E^{c})\cdot 0.\qed$
+Using Proposition 5.1 and Lemma 6.3, our main result in this subsection bounds the bias of the $d$-times iterated derivative of $f\in\operatorname{Poly}_{d+1}$ in terms of (polarizations of) the components of $f$.
 Proposition 6.4.
-For $f\in\operatorname{Poly}_{d+1}(\mathbb{F})$ where $\operatorname{ch}(\mathbb{F})\neq 2$, with $g=f_{d+1}$ and $h=f_{d}$,
+For $f\in\operatorname{Poly}_{d+1}(\mathbb{F})$ where $\operatorname{ch}(\mathbb{F})>2$, with $g=f_{d+1}$ and $h=f_{d}$,
 $\operatorname{bias}(\Delta^{d}f)\leq\operatorname*{\mathbb{E}}_{c}\operatorname{bias}(\overline{h}-\partial_{c}\overline{g})\quad\text{ and }\quad\operatorname{bias}(\Delta^{d}f)\leq\operatorname{bias}(\overline{g}).$
@@
 Thus, $\Delta^{d}_{\mathbf{v}}f(x)=\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v})$, and therefore
-$\operatorname{\mathfrak{bias}}(\Delta^{d}_{\mathbf{v}}f(x))=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v}))=\operatorname*{\mathbb{E}}_{x^{\prime}}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v}))=\operatorname{\mathfrak{bias}}(\overline{g}+\overline{h}).$
+$\operatorname{\mathfrak{bias}}(\Delta^{d}f)=\operatorname{\mathfrak{bias}}(\Delta^{d}_{\mathbf{v}}f(x))=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v}))=\operatorname*{\mathbb{E}}_{x^{\prime}}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v}))=\operatorname{\mathfrak{bias}}(\overline{g}+\overline{h}).$
 This has two implications. First, it follows from Lemma 6.3 that
@@
 $=\operatorname*{\mathbb{E}}_{c}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\partial_{-c}\,\overline{g}(\mathbf{v})+\overline{h}(\mathbf{v}))=\operatorname*{\mathbb{E}}_{c}\operatorname{\mathfrak{bias}}(\overline{h}-\partial_{c}\overline{g}).$
-These complete the proof. ∎
+Taking absolute value and using the triangle inequality completes the proof. ∎
 6.2 Main proof
-We now put everything together and prove Theorem 1.
+We now put everything together—in particular, Propositions 6.4, 4.5, 4.7 and Theorem 2.1—to prove Theorem 1.
 Theorem (1).
@@
 Let $g=f_{d+1}$ and $h=f_{d}$ be the components of $f$ of degrees $d+1$ and $d$, respectively. We first claim that, for any direction vector $c$,
-$\operatorname{rk}(g)\leq\operatorname{rk}(\overline{g})\quad\text{ and }\quad\operatorname{rk}(h-\partial_{c}g)\leq\operatorname{rk}(\overline{h}-\partial_{c}\overline{g}).$ (7)
+$\operatorname{rk}(h-\partial_{c}g)\leq\operatorname{rk}(\overline{h}-\partial_{c}\overline{g})\quad\text{ and }\quad\operatorname{rk}(g)\leq\operatorname{rk}(\overline{g}).$ (7)
 To see this, abbreviate $\mathbf{x}^{t}=(x,\ldots,x)$ ($t$ times), and observe that $\partial_{c}\overline{g}(\mathbf{x}^{d})=\overline{\partial_{c}g}(\mathbf{x}^{d})=d!\partial_{c}g(x)$, using Lemma 5.2 and Fact 5.7.
 Moreover by Fact 5.7, we have $\overline{h}(\mathbf{x}^{d})=d!h(x)$ and $\overline{g}(\mathbf{x}^{d+1})=(d+1)!g(x)$.
-As $\operatorname{ch}(\mathbb{F})>d+1$, we thus have $g(x)=\overline{g}(\mathbf{x}^{d})/(d+1)!$ and $(h-\partial_{c}g)(x)=(\overline{h}-\partial_{c}\overline{g})(\mathbf{x}^{d})/d!$, proving (7).
+As $\operatorname{ch}(\mathbb{F})>d+1$, we thus have $g(x)=\overline{g}(\mathbf{x}^{d+1})/(d+1)!$ and $(h-\partial_{c}g)(x)=(\overline{h}-\partial_{c}\overline{g})(\mathbf{x}^{d})/d!$, proving (7) (since plugging in the diagonal in a reducible multilinear form gives a reducible form).
 By Proposition 6.4, $\operatorname{bias}(\Delta^{d}f)\leq\operatorname{bias}(\overline{g})$, and $\operatorname{bias}(\Delta^{d}f)\leq\operatorname{bias}(\overline{h}-\partial_{c}\overline{g})$ for some direction vector $c$ which we henceforth fix.
-Put $\operatorname{bias}(\Delta^{d}f)=q^{-r}$, so that
+Put $\operatorname{bias}(\Delta^{d}f)=q^{-r}$ (if $\operatorname{bias}(\Delta^{d}f)=0$ there is nothing to prove), so that
 $\operatorname{bias}(\overline{g})\geq q^{-r}\quad\text{ and }\quad\operatorname{bias}(\overline{h}-\partial_{c}\overline{g})\geq q^{-r}.$ (8)
-Now, combine (7), (8), and Theorem 2.1 to deduce that
+Now, combine (7), (8), and Theorem 2.1 (using $\operatorname{rk}(T)\leq\operatorname{PR}(T)$) to deduce that
 $\operatorname{rk}(g)\leq\operatorname{rk}(\overline{g})\leq\tilde{O}_{d}(r)\quad\text{ and }\quad\operatorname{rk}(h-\partial_{c}g)\leq\operatorname{rk}(\overline{h}-\partial_{c}\overline{g})\leq\tilde{O}_{d}(r).$
 Apply Proposition 4.5 to obtain
-$\operatorname{rk}^{*}(g+h)\leq 2\operatorname{rk}(g)+\operatorname{rk}(h-\partial_{c}g)\leq\tilde{O}_{d}(r).$
-Next, apply Proposition 4.7 with $k=d+1$ ($\geq 3$ by assumption) on $g+h\in\operatorname{Poly}_{d+1}$. We deduce that there exists a polynomial $P$ with $\deg(P)\leq k-2=d-1$ such that
-$\operatorname{cor}(g+h,\,P)\geq q^{-2\operatorname{rk}^{*}(g+h)}\geq 1/q^{\tilde{O}_{d}(r)}.$
-To finish, write $f=g+h+f_{<d-1}$. We claim that $P+f_{<d-1}\in\operatorname{Poly}_{d-1}$ significantly correlates with $f$; indeed,
-$\operatorname{cor}_{<d}(f)$ $\geq\operatorname{cor}(f,\,P+f_{<d-1})=\operatorname{bias}(f-(P+f_{<d-1}))=\operatorname{bias}(g+h-P)$
-$=\operatorname{cor}(g+h,\,P)\geq 1/q^{\tilde{O}_{d}(r)}=\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}.$
+$\operatorname{rk}^{*}(g+h+\gamma)\leq\operatorname{rk}(g)+\operatorname{rk}(h-\partial_{c}g)\leq\tilde{O}_{d}(r)$
+for some $\gamma\in\operatorname{Poly}_{d-1}$.
+Next, apply Proposition 4.7 with $k=d+1$ ($\geq 3$ by assumption) on $g+h+\gamma\in\operatorname{Poly}_{d+1}$.
+We deduce that there exists a polynomial $P$ with $\deg(P)\leq k-2=d-1$ such that
+$\operatorname{cor}(g+h+\gamma,\,P)\geq q^{-2\operatorname{rk}^{*}(g+h+\gamma)}\geq q^{-\tilde{O}_{d}(r)}.$
+To finish, write $f=g+h+f_{<d}$. Note that $P+f_{<d}-\gamma\in\operatorname{Poly}_{d-1}$; we claim that it significantly correlates with $f$. Indeed,
+$\operatorname{cor}_{<d}(f)$ $\geq\operatorname{cor}(f,\,P+f_{<d}-\gamma)=\operatorname{bias}(f-(P+f_{<d}-\gamma))=\operatorname{bias}(g+h+\gamma-P)$
+$=\operatorname{cor}(g+h+\gamma,\,P)\geq q^{-\tilde{O}_{d}(r)}=\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}.$
 This completes the proof. ∎